### v1

# Flow-GRPO: Training Flow Matching Models via Online RL

Abstract
We propose Flow-GRPO, the first method integrating online reinforcement learning (RL) into flow matching models. Our approach uses two key strategies: (1) an ODE-to-SDE conversion that transforms a deterministic Ordinary Differential Equation (ODE) into an equivalent Stochastic Differential Equation (SDE) that matches the original model’s marginal distribution at all timesteps, enabling statistical sampling for RL exploration; and (2) a Denoising Reduction strategy that reduces training denoising steps while retaining the original inference timestep number, significantly improving sampling efficiency without performance degradation. Empirically, Flow-GRPO is effective across multiple text-to-image tasks. For complex compositions, RL-tuned SD3.5 generates nearly perfect object counts, spatial relations, and fine-grained attributes, boosting GenEval accuracy from $63\%$ to $95\%$. In visual text rendering, its accuracy improves from $59\%$ to $92\%$, significantly enhancing text generation. Flow-GRPO also achieves substantial gains in human preference alignment. Notably, little to no reward hacking occurred, meaning rewards did not increase at the cost of image quality or diversity, and both remained stable in our experiments.
Figure 1: (a) GenEval performance rises steadily throughout Flow-GRPO’s training and outperforms GPT-4o. (b) Image quality metrics on DrawBench [1] remain essentially unchanged. (c) Human Preference Scores on DrawBench improves after training. Results show that Flow-GRPO boosts the desired capability without degrading image quality or exhibiting reward-hacking behaviour.
1 Introduction
Flow matching [2, 3] models have become dominant in image generation [4, 5] due to their solid theoretical foundations and strong performance in producing high-quality images. However, they often struggle with composing complex scenes involving multiple objects, attributes, and relationships [6, 7], as well as text rendering [8]. At the same time, online reinforcement learning (RL) [9] has proven highly effective in enhancing the reasoning capabilities of large language models (LLMs) [10, 11]. While previous research has mainly focused on applying RL to early diffusion-based generative models [12] and offline RL techniques like direct preference optimization [13] for flow-based generative models [14, 15], the potential of online RL in advancing flow matching generative models remains largely unexplored. In this study, we explore how online RL can be leveraged to effectively improve flow matching models.
Training flow models with RL presents several critical challenges: (1) Flow models rely on a deterministic generative process based on ordinary differential equations (ODEs) [3], meaning they cannot sample stochastically during inference. In contrast, RL relies on stochastic sampling to explore environments, where the agent learns by trying different actions and receiving feedback. This need for stochasticity in RL conflicts with the deterministic nature of flow matching models. (2) Online RL requires efficient sampling to gather training data, but flow models typically involve many iterative steps to generate each sample, significantly reducing sampling efficiency. This issue becomes even more critical with large, advanced models [5, 4], which are computationally intensive. Therefore, to make RL effective in tasks such as image or video generation, improving sampling efficiency within RL training is crucial.
To address these challenges, we propose Flow-GRPO, which integrates GRPO [16] into flow matching models for text-to-image (T2I) generation, using two key strategies. First, we adopt the ODE-to-SDE strategy to overcome the deterministic nature of the original flow model. By converting the ODE-based flow into an equivalent Stochastic Differential Equation (SDE) framework, we introduce randomness while preserving the original marginal distributions. Second, to improve sampling efficiency in online RL, we apply the Denoising Reduction strategy—reducing denoising steps during training while keeping the full schedule during inference. Our experiments show that using fewer steps maintains performance while significantly reducing data generation costs.
We evaluate Flow-GRPO on T2I tasks with various reward types. (1) Verifiable rewards, using the GenEval [17] benchmark and visual text rendering task. GenEval includes compositional image generation tasks (e.g., generating specific object counts, colors, and spatial relationships), which can be automatically assessed with traditional detection methods. Flow-GRPO improves the accuracy of Stable Diffusion 3.5 Medium (SD3.5-M) [4] from 63% to 95% on GenEval, outperforming the state-of-the-art GPT-4o [18] model. For visual text rendering, SD3-M’s accuracy increases from 59% to 92%, greatly enhancing its text generation ability. (2) Model-based rewards, such as the human preference Pickscore [19] reward, show that our framework is task independent, demonstrating its generalizability and robustness. Importantly, all improvements are achieved with very little reward hacking, as demonstrated in Figure 1.
To summarize, the contributions of Flow-GRPO are as follows:
  • We are the first to introduce GRPO to flow matching models by converting deterministic ODE sampling into SDE sampling, showing that online RL is highly effective for T2I tasks. Specifically, our Flow-GRPO improves SD3.5-Medium accuracy from 63% to 95% without compromising image quality.
  • We find that online RL for flow matching models can use fewer denoising steps for training sample generation, significantly accelerating training. Full denoising steps are still used during testing to maintain performance.
  • We show that the Kullback-Leibler (KL) constraint is effective in preventing reward hacking, where reward increases but image quality or diversity decreases. We also emphasize that KL regularization is not empirically equivalent to early stopping. By adding an appropriate KL, we can achieve the same high reward as the KL-free version while maintaining image quality, though with a longer training time.
2 Related Work
RL for LLM.
Online RL has proven effective in boosting the reasoning abilities of large language models (LLMs), notably DeepSeek-R1 [10] and OpenAI-o1 [11]. By iteratively alternating between exploration and exploitation, these systems generate increasingly detailed responses and steadily enhance overall performance. Most successful pipelines employ policy gradient algorithms such as Proximal Policy Optimization (PPO) [20] and the value-free Group Relative Policy Optimization (GRPO) [16]. Because GRPO dispenses with a separate value network, it is more memory efficient; hence we adopt GRPO in this work. Nonetheless, PPO could be incorporated into a flow matching model through an analogous procedure.
Diffusion and Flow Matching.
Diffusion-based [21, 22, 23] generative models create a forward Markov chain that incrementally adds Gaussian noise to data and train a neural network to approximate the reverse-time denoising vector field; integrating this learned field by discrete DDPM steps or probability flow SDE solvers can yields high fidelity samples. Flow matching [2, 3] instead treats generation as learning a continuous-time normalizing flow whose vector field is trained by directly matching the velocity, enables deterministic sampling in a few ODE steps. Thanks to its efficiency, flow matching attains competitive FID with orders-of-magnitude fewer denoising steps than diffusion samplers and has already become the dominant approach in current image generation [4, 5] and video generation [24, 25, 26, 27] models.  [28, 29] unifies flow-based and diffusion-based methods using the SDE and ODE perspective. Our work builds on their theoretical foundations and introduce GRPO to flow-based models.
Alignment for T2I.
Recent work on aligning pretrained T2I models with human preferences has pursued several complementary directions. The main approaches are: (1) direct fine-tuning with scalar reward signals [30, 31, 32, 33]; (2) Reward-Weighted Regression (RWR) [34, 35, 36]; (3) Direct Preference Optimization (DPO) and related variants,[37, 38, 39, 14, 40, 41, 42, 43, 44, 36]; (4) Proximal Policy Optimization (PPO)-style policy gradients [45, 46, 47, 48, 49]; and (5) training-free alignment techniques [50, 51, 52]. These methods have successfully aligned T2I models with human preferences, improving aesthetics and semantic consistency. Building on this progress, we introduce GRPO for flow-matching models—the backbone of today’s state-of-the-art T2I systems. Concurrent work [53] apply GRPO to text-to-speech flow models, but instead of converting the ODE to an SDE to inject stochasticity, they reformulate velocity prediction as estimating a Gaussian distribution (predicting both the mean and variance of velocity), which necessitates retraining the pretrained model. [54] also recognize that stochasticity can be introduced from an SDE perspective, yet their work concentrates on inference-time scaling.
3 Preliminaries
In this section, we introduce the mathematical formulation of flow matching (Section 3.1) and describe how the denoising process can be mapped as a multi-step MDP (Section 3.2).
3.1 Flow Matching
Let ${\bm{x}}_{0}\sim X_{0}$ denote a data sample drawn from the true data distribution, and ${\bm{x}}_{1}\sim X_{1}$ denote a noise sample. Recent advanced image-generation models (e.g., [4, 5]) and video-generation models (e.g., [24, 26, 25, 27]) adopt the Rectified Flow [3] framework, which defines the “noised” data ${\bm{x}}_{t}$ as
${\bm{x}}_{t}=(1-t)\,{\bm{x}}_{0}\;+\;t\,{\bm{x}}_{1},$ (1)
for $t\in[0,1]$. Then a transformer model are trained to directly regress the velocity field ${\bm{v}}_{\theta}({\bm{x}}_{t},t)$ by minimizing the Flow Matching objective [2, 3]:
$\mathcal{L}(\theta)=\mathbb{E}_{t,\;{\bm{x}}_{0}\sim X_{0},\;{\bm{x}}_{1}\sim X_{1}}\bigl{[}\;\|{\bm{v}}\;-\;{\bm{v}}_{\theta}({\bm{x}}_{t},t)\|^{2}\bigr{]},$ (2)
where the target velocity field is ${\bm{v}}={\bm{x}}_{1}-{\bm{x}}_{0}$.
Figure 2: Overview of Flow-GRPO. Given a prompt set, we introduce an ODE-to-SDE strategy to enable stochastic sampling for online RL. With Denoising Reduction (only T = 10 steps), we efficiently gather low-quality but still informative trajectories. Rewards from these trajectories feed the GRPO loss, which updates the model online and yields an aligned policy.
3.2 Denoising as an MDP
According to [12], the iterative denoising process in diffusion models can be framed as an Markov decision process (MDP) $(\mathcal{S},\mathcal{A},\rho_{0},P,R)$. The state at time step p $t$ is ${\bm{s}}_{t}\!\triangleq\!({\bm{c}},t,{\bm{x}}_{t})$, the action is the denoised sample ${\bm{a}}_{t}\!\triangleq\!{\bm{x}}_{t-1}$ produced by the network, and the policy is the conditional model $\pi({\bm{a}}_{t}\mid{\bm{s}}_{t})\!\triangleq\!p_{\theta}({\bm{x}}_{t-1}\mid{\bm{x}}_{t},{\bm{c}})$. The transition kernel is deterministic $P({\bm{s}}_{t+1}\mid{\bm{s}}_{t},{\bm{a}}_{t})\!\triangleq\!(\delta_{\bm{c}},\delta_{t-1},\delta_{{\bm{x}}_{t-1}})$ and the initial-state distribution is $\rho_{0}({\bm{s}}_{0})\!\triangleq\!(p({\bm{c}}),\delta_{T},\mathcal{N}(\mathbf{0},\mathbf{I}))$, in which $\delta_{y}$ is the Dirac delta distribution with nonzero density only at $y$. The reward is non-zero only at the final step, $R({\bm{s}}_{t},{\bm{a}}_{t})\!\triangleq\!r({\bm{x}}_{0},{\bm{c}})$ if $t\!=\!0$ and 00 otherwise.
4 Flow-GRPO
In this section, we present the Flow-GRPO method, a GRPO-inspired algorithm that improves flow models through online RL. We begin by revisiting the core concept of GRPO [16] and recasting it in the context of flow matching. Next, we discuss how to convert the deterministic probability-flow ODE sampler into a reverse-time SDE with equivalent marginal distribution, thereby introducing stochasticity into the generation process—an essential ingredient for applying GRPO or other RL techniques. Finally, we introduce Denoise Reduction (Section 4.3), a practical data sampling method that dramatically accelerates training without degrading performance.
4.1 GRPO on Flow Matching
Reinforcement learning aims to learn a policy that maximizes expected cumulative reward under a given environment. Typically the goal is to optimize a policy $\pi_{\theta}$ with a regularized objective of the form:
$\max_{\theta}\mathbb{E}_{({\bm{s}}_{0},{\bm{a}}_{0},\dots,{\bm{s}}_{T},{\bm{a}}_{T})\sim\pi_{\theta}}\Bigg{[}\sum_{t=0}^{T}\Bigg{(}R({\bm{s}}_{t},{\bm{a}}_{t})-\beta D_{\text{KL}}(\pi_{\theta}(\cdot\mid{\bm{s}}_{t})||\pi_{\text{ref}}(\cdot\mid{\bm{s}}_{t}))\Bigg{)}\Bigg{]}.$ (3)
Among various RL algorithms, GRPO [16] provides a lightweight alternative to policy based methods such as PPO [20]. Unlike PPO [20], which requires training a separate value function—normally a separate model of similar size to the policy network, GRPO introduces a group relative formulation to estimate the advantage. In our setting, we recast the GRPO framework into the context of flow matching. Recall that the denoising process can be formulated as an MDP, as shown in Section 3.2. Given a prompt ${\bm{c}}$, the flow model $p_{\theta}$ samples a group of $G$ individual images $\{{\bm{x}}^{i}_{0}\}_{i=1}^{G}$ and the corresponding reverse-time trajectories $\{({\bm{x}}^{i}_{T},{\bm{x}}^{i}_{T-1},\cdots,{\bm{x}}^{i}_{0})\}_{i=1}^{G}$. Then, the advantage of the $i$-th image is calculated by normalizing the group-level rewards as follows:
$\hat{A}^{i}_{t}=\frac{R({\bm{x}}^{i}_{0},{\bm{c}})-\text{mean}(\{R({\bm{x}}^{i}_{0},{\bm{c}})\}_{i=1}^{G})}{\text{std}(\{R({\bm{x}}^{i}_{0},{\bm{c}})\}_{i=1}^{G})}.$ (4)
GRPO employs a clipped objective of Eq. 3, together with a KL penalty term:
$\mathcal{J}_{\text{Flow-GRPO}}(\theta)$ $=\mathbb{E}_{{\bm{c}}\sim\mathcal{C},\{{\bm{x}}^{i}\}_{i=1}^{G}\sim\pi_{\theta_{\text{old}}}(\cdot\mid{\bm{c}})}$ (5)
$\Bigg{[}\frac{1}{G}\sum_{i=1}^{G}\frac{1}{T}\sum_{t=0}^{T-1}\Bigg{(}\min\Big{(}r^{i}_{t}(\theta)\hat{A}^{i}_{t},\ \text{clip}\Big{(}r^{i}_{t}(\theta),1-\varepsilon,1+\varepsilon\Big{)}\hat{A}^{i}_{t}\Big{)}-\beta D_{\text{KL}}(\pi_{\theta}||\pi_{\text{ref}})\Bigg{)}\Bigg{]},$
where
$r^{i}_{t}(\theta)=\frac{p_{\theta}({\bm{x}}^{i}_{t-1}\mid{\bm{x}}^{i}_{t},{\bm{c}})}{p_{\theta_{\text{old}}}({\bm{x}}^{i}_{t-1}\mid{\bm{x}}^{i}_{t},{\bm{c}})}.$ (6)
4.2 ODE-to-SDE
In Eq. 4 and Eq. 6, GRPO relies on a stochastic sampling procedure in order to generate a diverse batch of trajectories for advantage estimation and policy exploration. In the case of diffusion models, the forward process consists of a sequence of random Gaussian noising steps, and its reverse-time sampler is implemented via a Markov chain of Gaussians with progressively decreasing variance—equivalently, an approximate solver for the score-based SDE. Thus, diffusion sampling is naturally stochastic.
While in the case of probability flow model, as previously discussed in Eq. 1, its forward process can be formulated as a deterministic ODE:
$\mathrm{d}{\bm{x}}_{t}={\bm{v}}_{t}\mathrm{d}t,$ (7)
where ${\bm{v}}_{t}$ is modeled via the flow matching objective in Eq. 2. A common sampling method is to discretize this ODE (e.g., via the explicit Euler method or a higher-order integrator) yielding a one-to-one mapping between successive time steps. For example, with a fixed step size $\Delta t$, the Euler scheme gives:
${\bm{x}}_{t-1}={\bm{x}}_{t}+\Delta t{\bm{v}}_{\theta}({\bm{x}}_{t},t,{\bm{c}})$ (8)
which produces a single, deterministic trajectory $\{{\bm{x}}_{T},{\bm{x}}_{T-1},\cdots,{\bm{x}}_{0}\}$ given any initial ${\bm{x}}_{T}$ and prompt ${\bm{c}}$.
This deterministic approach fails to meet GRPO’s policy update requirements in two key ways: (1) Eq. 6 requires computing $p({\bm{x}}_{t-1}\mid{\bm{x}}_{t},{\bm{c}})$, but deterministic sampling cannot provide probabilities. This limitation affects not only GRPO but all policy gradient methods, including the basic online REINFORCE [55] policy gradient, which also require probability computation. (2) Exploration is crucial for RL performance. As shown in Section 5.3, reduced randomness significantly lowers training efficiency. Deterministic sampling, which lacks randomness beyond the initial seed, is especially problematic.
To address this limitation, we convert the deterministic Flow-ODE from Eq. 7 into an equivalent SDE that matches the original model’s marginal probability density function at all timesteps. We outline the key process here. A detailed proof is provided in Appendix A. Following [23, 28, 29], we construct a reverse-SDE formulation that preserves the marginal distribution:
$\mathrm{d}{\bm{x}}_{t}=\left({\bm{v}}_{t}({\bm{x}}_{t})-\frac{\sigma_{t}^{2}}{2}\nabla\log p_{t}({\bm{x}}_{t})\right)\mathrm{d}t+\sigma_{t}\mathrm{d}{\bm{w}},$ (9)
where $\mathrm{d}{\bm{w}}$ denotes Wiener process increments and $\sigma_{t}$ control the level of stachasticity during generation. For rectified flow, the marginal score has a relation with velocity:
$\nabla\log p_{t}({\bm{x}})=-\frac{{\bm{x}}}{t}-\frac{1-t}{t}{\bm{v}}_{t}({\bm{x}}).$ (10)
Substituting Eq. 10 into Eq. 9 gives the final SDE:
$\mathrm{d}{\bm{x}}_{t}=\left[{\bm{v}}_{t}({\bm{x}}_{t})+\frac{\sigma_{t}^{2}}{2t}\left({\bm{x}}_{t}+(1-t){\bm{v}}_{t}({\bm{x}}_{t})\right)\right]\mathrm{d}t+\sigma_{t}\mathrm{d}{\bm{w}}.$ (11)
Applying Euler-Maruyama discretization yields the final update rule:
${\bm{x}}_{t+\Delta t}={\bm{x}}_{t}+\left[{\bm{v}}_{\theta}({\bm{x}}_{t},t)+\frac{\sigma_{t}^{2}}{2t}\big{(}{\bm{x}}_{t}+(1-t){\bm{v}}_{\theta}({\bm{x}}_{t},t)\big{)}\right]\Delta t+\sigma_{t}\sqrt{\Delta t}\,\epsilon$ (12)
where $\epsilon\sim\mathcal{N}(0,{\bm{I}})$ injects stochasticity. We use $\sigma_{t}=a\sqrt{\frac{t}{1-t}}$ in this paper, where $a$ is a scalar hyper-parameter that controls the noise level (See Section 5.3 for its impact on performance).
Eq. 12 reveals that the policy $\pi_{\theta}({\bm{x}}_{t-1}\mid{\bm{x}}_{t},{\bm{c}})$ is an isotropic Gaussian distribution. We can easily compute the KL divergence between $\pi_{\theta}$ and the reference policy $\pi_{\text{ref}}$ as a closed form:
$D_{\text{KL}}(\pi_{\theta}||\pi_{\text{ref}})=\frac{\|\overline{{\bm{x}}}_{t+\Delta t,\theta}-\overline{{\bm{x}}}_{t+\Delta t,\text{ref}}\|^{2}}{2\sigma_{t}^{2}\Delta t}=\frac{\Delta t}{2}\left(\frac{\sigma_{t}(1-t)}{2t}+\frac{1}{\sigma_{t}}\right)^{2}\|{\bm{v}}_{\theta}({\bm{x}}_{t},t)-{\bm{v}}_{\text{ref}}({\bm{x}}_{t},t)\|^{2}$ (13)
4.3 Denoising Reduction
To produce high-quality images, flow models typically require many denoising steps, making data collection costly for online RL. However, we find that large timesteps are unnecessary during online RL training. We can use significantly fewer denoising steps during sample generation, while retaining the original denoising steps during inference to get high-quality samples. Note that we set the timestep $T$ as $10$ in training, while the inference timestep $T$ is set as the original default setting ($T=40$) for SD3.5-M. Our experiments reveals that this approach enables fast training without sacrificing image quality at test time.
5 Experiments
In this section, we empirically evaluate Flow-GRPO’s ability to improving flow matching models on three representative tasks. (1) Composition Image Generation: This task requires precise object arrangement and attribute control. We report the results on GenEval. (2) Visual Text Rendering: a rule-based task that stresses the accurate rendering of the text specified in the prompt. (3) Human Preference Alignment: This task aims to align T2I models with human preferences. We use PickScore [19] as our reward model.
5.1 Experimental Setup
We introduce three tasks, detailing their respective prompts and reward definitions. For hyperparameter details and compute resource specifications, please refer to Appendix B.2 and Appendix B.2.1, respectively.
Compositional Image Generation
[…]
  • Counting: $r=|N_{\text{gen}}-N_{\text{ref}}|/N_{\text{ref}}$.
  • Position / Color: If the object count is correct, a partial reward is assigned; the remainder is granted when the predicted position or color is also correct.
Table 1: GenEval Result. Best scores are in blue, second-best in green. Results for models other than SD3.5-M are from [7] or their original papers. Obj.: Object; Attr.: Attribution.
Model Overall Single Obj. Two Obj. Counting Colors Position Attr. Binding
Diffusion Models
LDM [56] 0.37 0.92 0.29 0.23 0.70 0.02 0.05
SD1.5 [56] 0.43 0.97 0.38 0.35 0.76 0.04 0.06
SD2.1 [56] 0.50 0.98 0.51 0.44 0.85 0.07 0.17
SD-XL [57] 0.55 0.98 0.74 0.39 0.85 0.15 0.23
DALLE-2 [58] 0.52 0.94 0.66 0.49 0.77 0.10 0.19
DALLE-3 [58] 0.67 0.96 0.87 0.47 0.83 0.43 0.45
Autoregressive Models
Show-o [59] 0.53 0.95 0.52 0.49 0.82 0.11 0.28
Emu3-Gen [60] 0.54 0.98 0.71 0.34 0.81 0.17 0.21
JanusFlow [61] 0.63 0.97 0.59 0.45 0.83 0.53 0.42
Janus-Pro-7B [62] 0.80 0.99 0.89 0.59 0.90 0.79 0.66
GPT-4o [18] 0.84 0.99 0.92 0.85 0.92 0.75 0.61
Flow Matching Models
FLUX.1 Dev [5] 0.66 0.98 0.81 0.74 0.79 0.22 0.45
SD3.5-L [4] 0.71 0.98 0.89 0.73 0.83 0.34 0.47
SANA-1.5 4.8B [63] 0.81 0.99 0.93 0.86 0.84 0.59 0.65
SD3.5-M [4] 0.63 0.98 0.78 0.50 0.81 0.24 0.52
SD3.5-M+Flow-GRPO 0.95 1.00 0.99 0.95 0.92 0.99 0.86
Figure 3: Qualitative Comparison on the GenEval Benchmark. Our approach demonstrates superior performance in Counting, Colors, Attribute Binding, and Position.
Table 2: Performance on Compositional Image Generation, Visual Text Rendering, and Human Preference benchmarks, evaluated by task performance, image quality, and preference scores. Acc: Accuracy; ImgRwd: ImageReward; UniRwd: UnifiedReward.
Model Task Metric Image Quality Preference Score
GenEval OCR Acc. PickScore Aesthetic DeQA ImgRwd PickScore UniRwd
SD3.5-M 0.63 0.59 21.72 5.39 4.07 0.87 22.34 3.33
Compositional Image Generation
Flow-GRPO (w/o KL) 0.95 4.93 2.77 0.44 21.16 2.94
Flow-GRPO (w/ KL) 0.95 5.25 4.01 1.03 22.37 3.51
Visual Text Rendering
Flow-GRPO (w/o KL) 0.93 5.13 3.66 0.58 21.79 3.15
Flow-GRPO (w/ KL) 0.92 5.32 4.06 0.95 22.44 3.42
Human Preference Alignment
Flow-GRPO (w/o KL) 23.41 6.15 4.16 1.24 3.57
Flow-GRPO (w/ KL) 23.31 5.92 4.22 1.28 3.66
Visual Text Rendering [8]
Text is common in images such as posters, book covers, and memes, so the ability to place accurate and coherent text inside the generated images is crucial for T2V models. In our settings, we define an OCR task, where each prompt follows the template ‘‘A sign that says ‘‘text’’. Specifically, the placeholder ‘‘text’’ is the exact string that should appear in the image. We use GPT to produce 20K training prompts and 1K test prompts. Following [64], we measure text fidelity with the reward
$r=\frac{|N_{\text{gen}}-N_{\text{ref}}|}{N_{\text{ref}}},$
where $N_{\text{gen}}$ is the number of characters detected in the generated image and $N_{\text{ref}}$ is the number of characters inside the quotation marks in the prompt. This reward also serves as our measure of text accuracy.
Human Preference Alignment [19]
This task aims to align T2I models with human preferences. We use PickScore  [^2] https://huggingface.co/yuvalkirstain/PickScore_v1 as our reward model, which is based on large-scale human annotated pairwise comparisons of images generated from the same prompt. For each image and prompt pair, PickScore provides an overall score that evaluates multiple criteria, such as the alignment of the image with the prompt and its visual quality.
Image Quality Evaluation Metric
Because the T2I model is trained to maximise a predefined reward signal in RL, it is susceptible to reward hacking—situations in which the numerical reward rises but the quality or diversity of the generated images declines. This study aims to make online RL truly effective for T2I generation without compromising quality or diversity. To detect potential reward hacking beyond task-specific accuracy, we evaluate four automatic image quality metrics:
  • Aesthetic score [65]: a CLIP-based linear regressor that predicts an image’s aesthetic score on a 1–10 scale.
  • DeQA score [66]: a multimodal large language model based image-quality assessment (IQA) model that quantifies how distortions, texture damage, and other low-level artefacts affect perceived quality.
  • ImageReward [32]: a general purpose T2I human preference reward model that captures text–image alignment, visual fidelity, and harmlessness.
  • UnifiedReward [67]: a recently proposed unified reward model for multimodal understanding and generation that currently achieves state-of-the-art performance on the human preference assessment leaderboard.
All metrics are evaluated on DrawBench [1], a comprehensive and challenging benchmark with diverse prompts for T2I models.
5.2 Main Results
Figure 1 and Table 1 show Flow-GRPO’s GenEval performance steadily improving during training, ultimately outperforming GPT-4o. This occurs while maintaining both image quality metrics and preference scores on DrawBench, a benchmark with diverse and comprehensive prompts for evaluating general model capabilities. Figure 3 offers qualitative comparisons. Beyond Compositional Image Generation, Table 2 details evaluations on Visual Text Rendering and Human Preference tasks. Flow-GRPO improved text rendering ability, again without decreasing image quality metrics and preference scores on DrawBench. For the Human Preference task, image quality did not decrease with or without KL regularization. However, we found that omitting KL caused a collapse in visual diversity—a form of reward hacking discussed further in Section 5.3. These results demonstrate that Flow-GRPO boosts desired capabilities without degrading image quality or visual diversity.
5.3 Analysis
Reward Hacking
We explored two approaches to mitigate reward hacking: (1) Multi-reward ensemble: We combined preference-based rewards (e.g., UnifiedReward, PickScore, ImageReward), image quality rewards (e.g., Aesthetic Score, DeQA) with task reward. However, this also led to local blurriness and reduced diversity. Due to the challenge of covering all aspects of generation with diverse reward models, we abandoned this approach. (2) KL constraint: We tuned the KL coefficient to maintain a small, constant KL divergence during training, keeping the model close to its pretrained weight. This allows for task-specific reward optimization without harming general performance. As shown in Table 2, removing the KL constraint boosts performance on Compositional Image Generation and Visual Text Rendering but causes a sharp decline in image quality and preference scores on DrawBench. In contrast, a properly tuned KL constraint preserves quality while maintaining similar gains on task-specific metrics. In the Human Preference Alignment task, dropping the KL constraint doesn’t hurt image quality—likely due to PickScore overlap with evaluation metrics—but causes a collapse in visual diversity. Outputs converge to a single style, with different seeds producing nearly identical results. KL regularization prevents this collapse and preserves diversity. See Figure 7 in the Appendix C.2 for learning curves with and without KL, and Figure 4 for examples.
Figure 4: Effect of KL Regularization. The KL penalty effectively suppresses reward hacking, preventing potential Quality Degradation (for GenEval and OCR) and Diversity Decline (for PickScore).
Effect of Denoising Reduction
Figure 5 (a) highlights Denoising Reduction’s significant impact on accelerating training. To explore how different timesteps affect optimization, these experiments were conducted without the KL constraint. Reducing data collection timesteps from $40$ to $10$ achieved over a 4×4$\times$4 × speedup across all three tasks, without impacting final reward. While further tests showed this reduction didn’t consistently improve speed (and occasionally slowed some tasks), we standardized on $10$ timesteps for subsequent experiments, prioritizing the maintained reward. For the other two tasks, learning curves of reward versus training time are presented in Figure 6 in the Appendix C.1.
Effect of Noise Level
Higher $\sigma_{t}$ in the SDE boosts image diversity and exploration, vital for RL training. We control this exploration with a noise level $a$ (Eq. 12). Ablation studies (Figure 5 (b)) show $a$’s impact on performance. Small $a$ (e.g., 0.1) limits exploration and slows reward improvement. Increasing $a$ (up to 0.7) enhances exploration and accelerates reward gains. Beyond this, further increases (e.g., $a=0.7\rightarrow 1.0$) yield no extra benefit, as exploration is sufficient. However, too much noise degrades image quality, leading to zero reward and failed training. We advise using the highest noise level that maintains good image quality.
(a). Effect of Denoising Reduction on GenEval
(b). Effect of Noise Level Ablation on the OCR
Figure 5: Ablation studies on our critical design choices. (a) Denoising Reduction: Fewer denoising steps accelerate convergence and yield similar performance. (b) Noise Level: Moderate noise level ($a=0.7$) maximises OCR accuracy, while too little noise hampers exploration.
Generalization Analysis
Flow-GRPO demonstrates strong generalization on unseen scenarios from GenEval (Table 4). Specifically, it captures object number, color, and spatial relations, generalizing well to unseen object classes. It also effectively controls object count, generalizing from training on $2-4$ objects to generate $5-6$ objects. Furthermore, Table 3 shows Flow-GRPO achieves significant gains on T2I-CompBench++. This comprehensive benchmark for open-world compositional T2I generation features object classes and relationships substantially different from our model’s GenEval-style training data.
Table 3: T2I-CompBench++ Result. This evaluation uses the same model presented in Table 1, which was trained on the GenEval-generated dataset. The best score is in blue.
Model Color Shape Texture 2D-Spatial 3D-Spatial Numeracy Non-Spatial
Janus-Pro-7B [62] 0.5145 0.3323 0.4069 0.1566 0.2753 0.4406 0.3137
EMU3 [60] 0.7913 0.5846 0.7422
FLUX.1 Dev [5] 0.7407 0.5718 0.6922 0.2863 0.3866 0.6185 0.3127
SD3.5-M 0.7994 0.5669 0.7338 0.2850 0.3739 0.5927 0.3146
SD3.5-M+Flow-GRPO 0.8379 0.6130 0.7236 0.5447 0.4471 0.6752 0.3195
Table 4: Flow-GRPO demonstrates strong generalization. Unseen Objects: Trained on 60 object classes, evaluated on 20 unseen classes. Unseen Counting: Trained to render 2, 3, or 4 objects, was evaluated on rendering 5 or 6.
Method Unseen Objects Unseen Counting
Overall Single Obj. Two Obj. Counting Colors Position Attr. Binding Counting
SD3.5-M 0.64 0.96 0.73 0.53 0.87 0.26 0.47 0.13
SD3.5-M+Flow-GRPO 0.90 1.00 0.94 0.86 0.97 0.84 0.77 0.48
Limitations & Future Work.
While this work focuses on T2I tasks, Flow-GRPO’s potential applicability to video generation [25, 27] raises several questions for future research: (1) Reward Design for Video: Defining effective reward models for video generation is crucial. Simple heuristics like object detectors or trackers can promote physical realism and temporal consistency, but more sophisticated models are needed. (2) Balancing Multiple Rewards: Effective video generation requires optimizing multiple objectives (e.g., realism, smoothness, coherence). Balancing these often conflicting signals is challenging and demands careful tuning. (3) Scalability: Video generation is significantly more resource-intensive than T2I. Scaling Flow-GRPO for video thus demands more efficient data collection and training pipelines.
Appendix A Mathematical Derivations for Stochastic Sampling using Flow Models
We present a detailed proof here. To compute $p_{\theta}({\bm{x}}_{t-1}\mid{\bm{x}}_{t},{\bm{c}})$ in Equation 6 during forward sampling, we adapt flow models to a stochastic differential equation (SDE). While flow models normally follow a deterministic ODE:
$\mathrm{d}{\bm{x}}_{t}={\bm{v}}_{t}\mathrm{d}t$ (14)
We consider its stochastic counterpart. Inspired by the derivation from SDE to its probability flow ODE in SGMs [23], we aim to construct a SDE with specific drift and diffusion coefficients so that its marginal distribution matches that of Eq. 14. We begin with the generic form of SDE:
$\mathrm{d}{\bm{x}}_{t}=f_{\text{SDE}}({\bm{x}}_{t},t)\mathrm{d}t+\sigma_{t}\mathrm{d}{\bm{w}},$ (15)
Its marginal probability density $p_{t}({\bm{x}})$ evolves according to the Fokker–Planck equation [68], i.e.,
$\partial_{t}p_{t}(x)=-\nabla\cdot[f_{\text{SDE}}({\bm{x}}_{t},t)p_{t}({\bm{x}})]+\frac{1}{2}\nabla^{2}[\sigma_{t}^{2}p_{t}({\bm{x}})]$ (16)
Similarly, the marginal probability density associated with Eq. 14 evolves:
$\partial_{t}p_{t}({\bm{x}})$ $=-\nabla\cdot[{\bm{v}}_{t}({\bm{x}}_{t},t)p_{t}({\bm{x}})]$ (17)
To ensure that the stochastic process shares the same marginal distribution as the ODE, we impose:
$-\nabla\cdot[f_{\text{SDE}}p_{t}({\bm{x}})]+\frac{1}{2}\nabla^{2}[\sigma_{t}^{2}p_{t}({\bm{x}})]=-\nabla\cdot[{\bm{v}}_{t}p_{t}({\bm{x}})]$ (18)
Observing that
$\nabla^{2}[\sigma_{t}^{2}p_{t}({\bm{x}})]$ $=\sigma_{t}^{2}\nabla^{2}p_{t}({\bm{x}})$ (19)
$=\sigma_{t}^{2}\nabla\cdot(\nabla p_{t}({\bm{x}}))$
$=\sigma_{t}^{2}\nabla\cdot(p_{t}({\bm{x}})\nabla\log p_{t}({\bm{x}}))$
Substituting Eq. 19 to Eq. 18, we arrive at the drift coefficients of the target SDE:
$f_{\text{SDE}}={\bm{v}}_{t}-\frac{1}{2}\sigma_{t}^{2}\nabla\log p_{t}({\bm{x}})$ (20)
Hence we can rewrite the reverse-time SDE in Eq. 15 as:
$\mathrm{d}{\bm{x}}_{t}=\left({\bm{v}}_{t}({\bm{x}}_{t})-\frac{\sigma_{t}^{2}}{2}\nabla\log p_{t}({\bm{x}}_{t})\right)\mathrm{d}t+\sigma_{t}\mathrm{d}{\bm{w}},$ (21)
where $d{\bm{w}}$ denotes Wiener process increments and $\sigma_{t}$ is the diffusion coefficient that controls the level of stachasticity during sampling.
Once the score $\nabla\log p_{t}({\bm{x}}_{t})$ is available, the stochastic process can be sampled directly. While the Flow-Matching framework, this score is tied to the velocity field ${\bm{v}}_{t}$.
Specifically, let $\dot{\alpha_{t}}\equiv\partial\alpha_{t}/\partial t$. All expectations are over ${\bm{x}}_{0}\sim X_{0}$ and ${\bm{x}}_{1}\sim\mathcal{N}(0,{\bm{I}})$, where $X_{0}$ is the data distribution.
For the linear interpolation ${\bm{x}}_{t}=\alpha_{t}{\bm{x}}_{0}+\beta_{t}{\bm{x}}_{1}$, we have:
$p_{t|0}({\bm{x}}_{t}|{\bm{x}}_{0})=\mathcal{N}\left({\bm{x}}_{t}\mid\alpha_{t}{\bm{x}}_{0},\beta_{t}^{2}{\bm{I}}\right),$ (22)
yielding the conditional score:
$\nabla\log p_{t|0}({\bm{x}}_{t}|{\bm{x}}_{0})=-\frac{{\bm{x}}_{t}-\alpha_{t}{\bm{x}}_{0}}{\beta_{t}^{2}}=-\frac{{\bm{x}}_{1}}{\beta_{t}}.$ (23)
The marginal score becomes:
$\nabla\log p_{t}({\bm{x}}_{t})$ $=\mathbb{E}\left[\nabla\log p_{t|0}({\bm{x}}_{t}|{\bm{x}}_{0})\mid{\bm{x}}_{t}\right]$
$=-\frac{1}{\beta_{t}}\mathbb{E}[{\bm{x}}_{1}\mid{\bm{x}}_{t}].$ (24)
For the velocity field ${\bm{v}}_{t}({\bm{x}}_{t})$, we derive:
$\begin{split}{\bm{v}}_{t}({\bm{x}})&=\mathbb{E}\left[\dot{\alpha_{t}}{\bm{x}}_{0}+\dot{\beta_{t}}{\bm{x}}_{1}\mid{\bm{x}}_{t}={\bm{x}}\right]\\ &=\dot{\alpha_{t}}\mathbb{E}[{\bm{x}}_{0}\mid{\bm{x}}_{t}={\bm{x}}]+\dot{\beta_{t}}\mathbb{E}[{\bm{x}}_{1}\mid{\bm{x}}_{t}={\bm{x}}]\\ &=\dot{\alpha_{t}}\mathbb{E}\left[\frac{{\bm{x}}_{t}-\beta_{t}{\bm{x}}_{1}}{\alpha_{t}}\mid{\bm{x}}_{t}={\bm{x}}\right]+\dot{\beta_{t}}\mathbb{E}[{\bm{x}}_{1}\mid{\bm{x}}_{t}={\bm{x}}]\\ &=\frac{\dot{\alpha_{t}}}{\alpha_{t}}{\bm{x}}-\frac{\dot{\alpha_{t}}\beta_{t}}{\alpha_{t}}\mathbb{E}[{\bm{x}}_{1}\mid{\bm{x}}_{t}={\bm{x}}]+\dot{\beta_{t}}\mathbb{E}[{\bm{x}}_{1}\mid{\bm{x}}_{t}={\bm{x}}]\\ &=\frac{\dot{\alpha_{t}}}{\alpha_{t}}{\bm{x}}-\left(\dot{\beta_{t}}\beta_{t}-\frac{\dot{\alpha_{t}}\beta_{t}^{2}}{\alpha_{t}}\right)\nabla\log p_{t}({\bm{x}}),\end{split}$ (25)
Substituting $\alpha_{t}=1-t$ and $\beta_{t}=t$ simplifies Equation 25 to:
${\bm{v}}_{t}({\bm{x}})=-\frac{{\bm{x}}}{1-t}-\frac{t}{1-t}\nabla\log p_{t}({\bm{x}}).$ (26)
Solving for the score yields:
$\nabla\log p_{t}({\bm{x}})=-\frac{{\bm{x}}}{t}-\frac{1-t}{t}{\bm{v}}_{t}({\bm{x}}).$ (27)
Substituting Equation 27 into 21 gives the final SDE:
$\mathrm{d}{\bm{x}}_{t}=\left[{\bm{v}}_{t}({\bm{x}}_{t})+\frac{\sigma_{t}^{2}}{2t}\left({\bm{x}}_{t}+(1-t){\bm{v}}_{t}({\bm{x}}_{t})\right)\right]\mathrm{d}t+\sigma_{t}\mathrm{d}{\bm{w}}.$ (28)
Applying Euler-Maruyama discretization yields the update rule:
${\bm{x}}_{t+\Delta t}={\bm{x}}_{t}+\left[{\bm{v}}_{\theta}({\bm{x}}_{t},t)+\frac{\sigma_{t}^{2}}{2t}\big{(}{\bm{x}}_{t}+(1-t){\bm{v}}_{\theta}({\bm{x}}_{t},t)\big{)}\right]\Delta t+\sigma_{t}\sqrt{\Delta t}\,\epsilon,$ (29)
where $\epsilon\sim\mathcal{N}(0,{\bm{I}})$ injects stochasticity.
Appendix B Further Details on the Experimental Setup
B.1 Model Specification
The following table lists the base model and the reward models and their corresponding links.
B.2 Hyperparameters Specification
Except for $\beta$, GRPO hyperparameters are fixed across tasks. We use a sampling timestep $T=10$ and an evaluation timestep $T=40$. Other settings include a group size $G=24$, an noise level $a=0.7$ and an image resolution of 512. The KL ratio $\beta$ is set to 0.004 for GenEval and Text Rendering, and 0.001 for Pickscore. We use Lora with $\alpha=64$ and $r=32$.
B.2.1 Compute Resources Specification
We train our model using 24 NVIDIA A800 GPUs. The learning curves in Appendix C.2 provide details on the specific GPU hours.
Appendix C Extended Experimental Results
C.1 Effect of Denoising Reduction
We show the extended Denoising Reduction ablations of Visual Text Rendering and Human Preference Alignment tasks in Figure 6.
(a). Visual Text Rendering
(b). Human Preference Alignment
Figure 6: Effect of Denoising Reduction
C.2 Learning Curves with or without KL
Figure 7 shows learning curves for three tasks, with and without KL. These results emphasize that KL regularization is not empirically equivalent to early stopping. Adding appropriate KL can achieve the same high reward as the KL-free version and maintain image quality, though it requires longer training.
(a). Compositional Image Generation
(b). Visual Text Rendering
(c). Human Preference Alignment
Figure 7: Learning Curves with and without KL. KL penalty slows early training yet effectively suppresses reward hacking.

### v2

@@
 Flow-GRPO also achieves substantial gains in human preference alignment.
-Notably, little to no reward hacking occurred, meaning rewards did not increase at the cost of image quality or diversity, and both remained stable in our experiments.
+Notably, very little reward hacking occurred, meaning rewards did not increase at the cost of appreciable image quality or diversity degradation.
 Figure 1: (a) GenEval performance rises steadily throughout Flow-GRPO’s training and outperforms GPT-4o.
@@
 (c) Human Preference Scores on DrawBench improves after training.
-Results show that Flow-GRPO boosts the desired capability without degrading image quality or exhibiting reward-hacking behaviour.
+Results show that Flow-GRPO enhances the desired capability while preserving image quality and exhibiting minimal reward-hacking.
 1 Introduction
@@
 5.1 Experimental Setup
-We introduce three tasks, detailing their respective prompts and reward definitions.
-For hyperparameter details and compute resource specifications, please refer to Appendix B.2 and Appendix B.2.1, respectively.
+We introduce three tasks, detailing their respective prompts and reward definitions. For hyperparameter details and compute resource specifications, please refer to Appendix B.2 and Appendix B.2.1.
 Compositional Image Generation
@@
 $r=\frac{|N_{\text{gen}}-N_{\text{ref}}|}{N_{\text{ref}}},$
-where $N_{\text{gen}}$ is the number of characters detected in the generated image and $N_{\text{ref}}$ is the number of characters inside the quotation marks in the prompt.
-This reward also serves as our measure of text accuracy.
+where $N_{\text{gen}}$ is the number of characters detected in the image and $N_{\text{ref}}$ is the number of characters inside the quotation marks in the prompt.
+This reward also serves as our metric of text accuracy.
 Human Preference Alignment [19]
@@
 To detect potential reward hacking beyond task-specific accuracy, we evaluate four automatic image quality metrics:
-  • Aesthetic score [65]: a CLIP-based linear regressor that predicts an image’s aesthetic score on a 1–10 scale.
+  • Aesthetic score [66]: a CLIP-based linear regressor that predicts an image’s aesthetic score.
   • DeQA score [67]: a multimodal large language model based image-quality assessment (IQA) model that quantifies how distortions, texture damage, and other low-level artefacts affect perceived quality.
@@
 Flow-GRPO improved text rendering ability, again without decreasing image quality metrics and preference scores on DrawBench.
-For the Human Preference task, image quality did not decrease with or without KL regularization.
+See Figures 8 and 9 in Appendix C.3 for related qualitative examples.
+For the Human Preference task, image quality did not decrease without KL regularization.
 However, we found that omitting KL caused a collapse in visual diversity—a form of reward hacking discussed further in Section 5.3.
-These results demonstrate that Flow-GRPO boosts desired capabilities without degrading image quality or visual diversity.
+These results demonstrate that Flow-GRPO boosts desired capabilities while causing very little degradation to image quality or visual diversity.
 5.3 Analysis
@@
 Figure 7: Learning Curves with and without KL. KL penalty slows early training yet effectively suppresses reward hacking.
+C.3 Qualitative Results
+In addition to the GenEval reward comparisons in Figure 3, Figures 8 and 9 show qualitative comparisons between SD-3.5-M and SD-3.5-M + Flow-GRPO trained with PickScore and OCR rewards, respectively.
+Figure 8: Qualitative comparison between the SD-3.5-M and SD-3.5-M + Flow-GRPO trained with PickScore reward.
+Figure 9: Qualitative comparison between the SD-3.5-M and SD-3.5-M + Flow-GRPO trained with OCR reward.

### v3

@@
 Abstract
-We propose Flow-GRPO, the first method integrating online reinforcement learning (RL) into flow matching models.
-Our approach uses two key strategies: (1) an ODE-to-SDE conversion that transforms a deterministic Ordinary Differential Equation (ODE) into an equivalent Stochastic Differential Equation (SDE) that matches the original model’s marginal distribution at all timesteps, enabling statistical sampling for RL exploration; and (2) a Denoising Reduction strategy that reduces training denoising steps while retaining the original inference timestep number, significantly improving sampling efficiency without performance degradation.
+We propose Flow-GRPO, the first method to integrate online reinforcement learning (RL) into flow matching models.
+Our approach uses two key strategies: (1) an ODE-to-SDE conversion that transforms a deterministic Ordinary Differential Equation (ODE) into an equivalent Stochastic Differential Equation (SDE) that matches the original model’s marginal distribution at all timesteps, enabling statistical sampling for RL exploration; and (2) a Denoising Reduction strategy that reduces training denoising steps while retaining the original number of inference steps, significantly improving sampling efficiency without sacrificing performance.
 Empirically, Flow-GRPO is effective across multiple text-to-image tasks.
-For complex compositions, RL-tuned SD3.5 generates nearly perfect object counts, spatial relations, and fine-grained attributes, boosting GenEval accuracy from $63\%$ to $95\%$.
-In visual text rendering, its accuracy improves from $59\%$ to $92\%$, significantly enhancing text generation.
+For compositional generation, RL-tuned SD3.5-M generates nearly perfect object counts, spatial relations, and fine-grained attributes, increasing GenEval accuracy from $63\%$ to $95\%$.
+In visual text rendering, accuracy improves from $59\%$ to $92\%$, greatly enhancing text generation.
 Flow-GRPO also achieves substantial gains in human preference alignment.
@@
 1 Introduction
-Flow matching [2, 3] models have become dominant in image generation [4, 5] due to their solid theoretical foundations and strong performance in producing high-quality images.
+Flow matching [2, 3] models have become dominant in image generation [4, 5] due to their solid theoretical foundations and strong performance in producing high quality images.
 However, they often struggle with composing complex scenes involving multiple objects, attributes, and relationships [6, 7], as well as text rendering [8].
@@
 In this study, we explore how online RL can be leveraged to effectively improve flow matching models.
-Training flow models with RL presents several critical challenges: (1) Flow models rely on a deterministic generative process based on ordinary differential equations (ODEs) [3], meaning they cannot sample stochastically during inference.
-In contrast, RL relies on stochastic sampling to explore environments, where the agent learns by trying different actions and receiving feedback.
+Training flow models with RL presents several critical challenges: (1) Flow models rely on a deterministic generative process based on ODEs [3], meaning they cannot sample stochastically during inference.
+In contrast, RL relies on stochastic sampling to explore the environment, learning by trying different actions and improving based on rewards.
 This need for stochasticity in RL conflicts with the deterministic nature of flow matching models.
-(2) Online RL requires efficient sampling to gather training data, but flow models typically involve many iterative steps to generate each sample, significantly reducing sampling efficiency.
-This issue becomes even more critical with large, advanced models [5, 4], which are computationally intensive.
-Therefore, to make RL effective in tasks such as image or video generation, improving sampling efficiency within RL training is crucial.
+(2) Online RL depends on efficient sampling to collect training data, but flow models typically require many iterative steps to generate each sample, limiting efficiency.
+This issue is more pronounced with large models [5, 4].
+To make RL practical for tasks like image or video generation, improving sampling efficiency is essential.
 To address these challenges, we propose Flow-GRPO, which integrates GRPO [16] into flow matching models for text-to-image (T2I) generation, using two key strategies.
@@
 By converting the ODE-based flow into an equivalent Stochastic Differential Equation (SDE) framework, we introduce randomness while preserving the original marginal distributions.
-Second, to improve sampling efficiency in online RL, we apply the Denoising Reduction strategy—reducing denoising steps during training while keeping the full schedule during inference.
+Second, to improve sampling efficiency in online RL, we apply the Denoising Reduction strategy, which reduces denoising steps during training while keeping the full schedule during inference.
 Our experiments show that using fewer steps maintains performance while significantly reducing data generation costs.
@@
 (1) Verifiable rewards, using the GenEval [17] benchmark and visual text rendering task.
-GenEval includes compositional image generation tasks (e.g., generating specific object counts, colors, and spatial relationships), which can be automatically assessed with traditional detection methods.
+GenEval includes compositional image generation tasks (e.g., generating specific object counts, colors, and spatial relationships), which can be automatically assessed with object detection methods.
 Flow-GRPO improves the accuracy of Stable Diffusion 3.5 Medium (SD3.5-M) [4] from 63% to 95% on GenEval, outperforming the state-of-the-art GPT-4o [18] model.
-For visual text rendering, SD3-M’s accuracy increases from 59% to 92%, greatly enhancing its text generation ability.
-(2) Model-based rewards, such as the human preference Pickscore [19] reward, show that our framework is task independent, demonstrating its generalizability and robustness.
+For visual text rendering, SD3.5-M’s accuracy increases from 59% to 92%, greatly enhancing its text generation ability.
+(2) Model-based rewards, such as the human preference Pickscore [19] reward.
+These results show that our framework is task independent, demonstrating its generalizability and robustness.
 Importantly, all improvements are achieved with very little reward hacking, as demonstrated in Figure 1.
 To summarize, the contributions of Flow-GRPO are as follows:
-  • We are the first to introduce GRPO to flow matching models by converting deterministic ODE sampling into SDE sampling, showing that online RL is highly effective for T2I tasks.
-Specifically, our Flow-GRPO improves SD3.5-Medium accuracy from 63% to 95% without compromising image quality.
-  • We find that online RL for flow matching models can use fewer denoising steps for training sample generation, significantly accelerating training.
-Full denoising steps are still used during testing to maintain performance.
-  • We show that the Kullback-Leibler (KL) constraint is effective in preventing reward hacking, where reward increases but image quality or diversity decreases.
-We also emphasize that KL regularization is not empirically equivalent to early stopping.
-By adding an appropriate KL, we can achieve the same high reward as the KL-free version while maintaining image quality, though with a longer training time.
+  • We are the first to introduce GRPO to flow matching models by converting deterministic ODE sampling into SDE sampling, showing the effectiveness of online RL for T2I tasks.
+Flow-GRPO improves SD3.5-M accuracy from 63% to 95% without noticeably compromising image quality.
+  • We find that online RL for flow matching models does not require the standard long timesteps for training sample collection.
+By using fewer denoising steps during training and retaining the original steps during testing, we can significantly accelerate the training process.
+  • We show that the Kullback-Leibler (KL) constraint effectively prevents reward hacking, where reward increases at the cost of image quality or diversity.
+KL regularization is not empirically equivalent to early stopping.
+With a proper KL term, we can match the high reward of the KL-free version while preserving image quality, albeit with longer training.
 2 Related Work
 RL for LLM.
-Online RL has proven effective in boosting the reasoning abilities of large language models (LLMs), notably DeepSeek-R1 [10] and OpenAI-o1 [11].
-By iteratively alternating between exploration and exploitation, these systems generate increasingly detailed responses and steadily enhance overall performance.
-Most successful pipelines employ policy gradient algorithms such as Proximal Policy Optimization (PPO) [20] and the value-free Group Relative Policy Optimization (GRPO) [16].
-Because GRPO dispenses with a separate value network, it is more memory efficient; hence we adopt GRPO in this work.
-Nonetheless, PPO could be incorporated into a flow matching model through an analogous procedure.
+Online RL has effectively improved the reasoning abilities of LLMs, such as DeepSeek-R1 [10] and OpenAI-o1 [11], using policy gradient methods like PPO [20] or value-free GRPO [16].
+GRPO is more memory efficient by removing the need for a value network, so we adopt it in this work.
+PPO can also be applied to flow matching in a similar way.
 Diffusion and Flow Matching.
-Diffusion-based [21, 22, 23] generative models create a forward Markov chain that incrementally adds Gaussian noise to data and train a neural network to approximate the reverse-time denoising vector field; integrating this learned field by discrete DDPM steps or probability flow SDE solvers can yields high fidelity samples.
-Flow matching [2, 3] instead treats generation as learning a continuous-time normalizing flow whose vector field is trained by directly matching the velocity, enables deterministic sampling in a few ODE steps.
-Thanks to its efficiency, flow matching attains competitive FID with orders-of-magnitude fewer denoising steps than diffusion samplers and has already become the dominant approach in current image generation [4, 5] and video generation [24, 25, 26, 27] models.
-[28, 29] unifies flow-based and diffusion-based methods using the SDE and ODE perspective.
-Our work builds on their theoretical foundations and introduce GRPO to flow-based models.
+Diffusion models [21, 22, 23] add Gaussian noise to data and train a neural network to reverse the process.
+Sampling uses discrete DDPM steps or probability flow SDE solvers to generate high-fidelity outputs.
+Flow matching [2, 3] learns a continuous-time normalizing flow by directly matching the velocity field, allowing efficient deterministic sampling with only a few ODE steps.
+It achieves competitive FID with far fewer denoising steps than diffusion, making it the dominant choice in recent image [4, 5] and video [24, 25, 26, 27] generation models.
+Recent work [28, 29] unifies diffusion and flow models under an SDE/ODE framework.
+Our work builds on their theoretical foundations and introduces GRPO to flow-based models.
 Alignment for T2I.
-Recent work on aligning pretrained T2I models with human preferences has pursued several complementary directions.
-The main approaches are: (1) direct fine-tuning with scalar reward signals [30, 31, 32, 33]; (2) Reward-Weighted Regression (RWR) [34, 35, 36]; (3) Direct Preference Optimization (DPO) and related variants,[37, 38, 39, 14, 40, 41, 42, 43, 44, 36]; (4) Proximal Policy Optimization (PPO)-style policy gradients [45, 46, 47, 48, 49]; and (5) training-free alignment techniques [50, 51, 52].
+Recent efforts to align pretrained T2I models with human preferences follow five main directions: (1) direct fine-tuning with scalar rewards [30, 31, 32, 33]; (2) Reward Weighted Regression (RWR) [34, 35, 36]; (3) Direct Preference Optimization (DPO) and variants [37, 38, 39, 14, 40, 41, 42, 43, 44, 36]; (4) PPO-style policy gradients [45, 46, 47, 48, 49, 50]; (5) training-free alignment methods [51, 52, 53].
 These methods have successfully aligned T2I models with human preferences, improving aesthetics and semantic consistency.
-Building on this progress, we introduce GRPO for flow-matching models—the backbone of today’s state-of-the-art T2I systems.
-Concurrent work [53] apply GRPO to text-to-speech flow models, but instead of converting the ODE to an SDE to inject stochasticity, they reformulate velocity prediction as estimating a Gaussian distribution (predicting both the mean and variance of velocity), which necessitates retraining the pretrained model.
-[54] also recognize that stochasticity can be introduced from an SDE perspective, yet their work concentrates on inference-time scaling.
+Building on this progress, we introduce GRPO for flow matching models, the backbone of today’s state-of-the-art T2I systems.
+Concurrent work [54] applies GRPO to text-to-speech flow models, but instead of converting the ODE to an SDE to inject stochasticity, they reformulate velocity prediction by estimating a Gaussian distribution (predicting both the mean and variance of velocity), which requires retraining the pre-trained model.
+Another study [55] also explores SDE-based stochasticity but focuses on inference-time scaling.
 3 Preliminaries
-In this section, we introduce the mathematical formulation of flow matching (Section 3.1) and describe how the denoising process can be mapped as a multi-step MDP (Section 3.2).
-3.1 Flow Matching
-Let ${\bm{x}}_{0}\sim X_{0}$ denote a data sample drawn from the true data distribution, and ${\bm{x}}_{1}\sim X_{1}$ denote a noise sample.
+In this section, we introduce the mathematical formulation of flow matching and describe how the denoising process can be mapped as a multi-step MDP.
+Flow Matching.
+Let ${\bm{x}}_{0}\sim X_{0}$ be a data sample from the true distribution, and ${\bm{x}}_{1}\sim X_{1}$ denote a noise sample.
 Recent advanced image-generation models (e.g., [4, 5]) and video-generation models (e.g., [24, 26, 25, 27]) adopt the Rectified Flow [3] framework, which defines the “noised” data ${\bm{x}}_{t}$ as
@@
 Rewards from these trajectories feed the GRPO loss, which updates the model online and yields an aligned policy.
-3.2 Denoising as an MDP
-According to [12], the iterative denoising process in diffusion models can be framed as an Markov decision process (MDP) $(\mathcal{S},\mathcal{A},\rho_{0},P,R)$.
-The state at time step p $t$ is ${\bm{s}}_{t}\!\triangleq\!({\bm{c}},t,{\bm{x}}_{t})$, the action is the denoised sample ${\bm{a}}_{t}\!\triangleq\!{\bm{x}}_{t-1}$ produced by the network, and the policy is the conditional model $\pi({\bm{a}}_{t}\mid{\bm{s}}_{t})\!\triangleq\!p_{\theta}({\bm{x}}_{t-1}\mid{\bm{x}}_{t},{\bm{c}})$.
-The transition kernel is deterministic $P({\bm{s}}_{t+1}\mid{\bm{s}}_{t},{\bm{a}}_{t})\!\triangleq\!(\delta_{\bm{c}},\delta_{t-1},\delta_{{\bm{x}}_{t-1}})$ and the initial-state distribution is $\rho_{0}({\bm{s}}_{0})\!\triangleq\!(p({\bm{c}}),\delta_{T},\mathcal{N}(\mathbf{0},\mathbf{I}))$, in which $\delta_{y}$ is the Dirac delta distribution with nonzero density only at $y$.
-The reward is non-zero only at the final step, $R({\bm{s}}_{t},{\bm{a}}_{t})\!\triangleq\!r({\bm{x}}_{0},{\bm{c}})$ if $t\!=\!0$ and 00 otherwise.
+Denoising as an MDP.
+As shown in [12], the iterative denoising process in flow matching models can be formulated as a Markov Decision Process (MDP) $(\mathcal{S},\mathcal{A},\rho_{0},P,R)$.
+The state at step $t$ is ${\bm{s}}_{t}\triangleq({\bm{c}},t,{\bm{x}}_{t})$, the action is the denoised sample ${\bm{a}}_{t}\triangleq{\bm{x}}_{t-1}$ predicted by the model, and the policy is $\pi({\bm{a}}_{t}\mid{\bm{s}}_{t})\triangleq p_{\theta}({\bm{x}}_{t-1}\mid{\bm{x}}_{t},{\bm{c}})$.
+The transition is deterministic: $P({\bm{s}}_{t+1}\mid{\bm{s}}_{t},{\bm{a}}_{t})\triangleq(\delta_{\bm{c}},\delta_{t-1},\delta_{{\bm{x}}_{t-1}})$, and the initial state distribution is $\rho_{0}({\bm{s}}_{0})\triangleq(p({\bm{c}}),\delta_{T},\mathcal{N}(\mathbf{0},\mathbf{I}))$, where $\delta_{y}$ is the Dirac delta distribution centered at $y$.
+The reward is only given at the final step: $R({\bm{s}}_{t},{\bm{a}}_{t})\triangleq r({\bm{x}}_{0},{\bm{c}})$ if $t=0$, and 0 otherwise.
 4 Flow-GRPO
-In this section, we present the Flow-GRPO method, a GRPO-inspired algorithm that improves flow models through online RL.
-We begin by revisiting the core concept of GRPO [16] and recasting it in the context of flow matching.
-Next, we discuss how to convert the deterministic probability-flow ODE sampler into a reverse-time SDE with equivalent marginal distribution, thereby introducing stochasticity into the generation process—an essential ingredient for applying GRPO or other RL techniques.
-Finally, we introduce Denoise Reduction (Section 4.3), a practical data sampling method that dramatically accelerates training without degrading performance.
-4.1 GRPO on Flow Matching
-Reinforcement learning aims to learn a policy that maximizes expected cumulative reward under a given environment.
-Typically the goal is to optimize a policy $\pi_{\theta}$ with a regularized objective of the form:
+In this section, we present Flow-GRPO, which enhances flow models using online RL.
+We begin by revisiting the core idea of GRPO [16] and adapting it to flow matching.
+We then show how to convert the deterministic ODE sampler into a SDE sampler with the same marginal distribution, introducing the stochasticity needed for applying GRPO.
+Finally, we introduce Denoise Reduction, a practical sampling strategy that significantly speeds up training without sacrificing performance.
+GRPO on Flow Matching.
+RL aims to learn a policy that maximizes the expected cumulative reward. This is often formulated as optimizing a policy $\pi_{\theta}$ with a regularized objective:
 $\max_{\theta}\mathbb{E}_{({\bm{s}}_{0},{\bm{a}}_{0},\dots,{\bm{s}}_{T},{\bm{a}}_{T})\sim\pi_{\theta}}\Bigg{[}\sum_{t=0}^{T}\Bigg{(}R({\bm{s}}_{t},{\bm{a}}_{t})-\beta D_{\text{KL}}(\pi_{\theta}(\cdot\mid{\bm{s}}_{t})||\pi_{\text{ref}}(\cdot\mid{\bm{s}}_{t}))\Bigg{)}\Bigg{]}.$ (3)
-Among various RL algorithms, GRPO [16] provides a lightweight alternative to policy based methods such as PPO [20].
-Unlike PPO [20], which requires training a separate value function—normally a separate model of similar size to the policy network, GRPO introduces a group relative formulation to estimate the advantage.
-In our setting, we recast the GRPO framework into the context of flow matching.
-Recall that the denoising process can be formulated as an MDP, as shown in Section 3.2.
+Unlike other policy based methods like PPO [20], GRPO [16] provides a lightweight alternative, which introduces a group relative formulation to estimate the advantage.
+Recall that the denoising process can be formulated as an MDP, as shown in Section 3.
 Given a prompt ${\bm{c}}$, the flow model $p_{\theta}$ samples a group of $G$ individual images $\{{\bm{x}}^{i}_{0}\}_{i=1}^{G}$ and the corresponding reverse-time trajectories $\{({\bm{x}}^{i}_{T},{\bm{x}}^{i}_{T-1},\cdots,{\bm{x}}^{i}_{0})\}_{i=1}^{G}$.
@@
 $\hat{A}^{i}_{t}=\frac{R({\bm{x}}^{i}_{0},{\bm{c}})-\text{mean}(\{R({\bm{x}}^{i}_{0},{\bm{c}})\}_{i=1}^{G})}{\text{std}(\{R({\bm{x}}^{i}_{0},{\bm{c}})\}_{i=1}^{G})}.$ (4)
-GRPO employs a clipped objective of Eq. 3, together with a KL penalty term:
-$\mathcal{J}_{\text{Flow-GRPO}}(\theta)$ $=\mathbb{E}_{{\bm{c}}\sim\mathcal{C},\{{\bm{x}}^{i}\}_{i=1}^{G}\sim\pi_{\theta_{\text{old}}}(\cdot\mid{\bm{c}})}$ (5)
-$\Bigg{[}\frac{1}{G}\sum_{i=1}^{G}\frac{1}{T}\sum_{t=0}^{T-1}\Bigg{(}\min\Big{(}r^{i}_{t}(\theta)\hat{A}^{i}_{t},\ \text{clip}\Big{(}r^{i}_{t}(\theta),1-\varepsilon,1+\varepsilon\Big{)}\hat{A}^{i}_{t}\Big{)}-\beta D_{\text{KL}}(\pi_{\theta}||\pi_{\text{ref}})\Bigg{)}\Bigg{]},$
+GRPO optimizes the policy model by maximizing the following objective:
+$\mathcal{J}_{\text{Flow-GRPO}}(\theta)=\mathbb{E}_{{\bm{c}}\sim\mathcal{C},\{{\bm{x}}^{i}\}_{i=1}^{G}\sim\pi_{\theta_{\text{old}}}(\cdot\mid{\bm{c}})}f(r,\hat{A},\theta,\varepsilon,\beta),$ (5)
 where
-$r^{i}_{t}(\theta)=\frac{p_{\theta}({\bm{x}}^{i}_{t-1}\mid{\bm{x}}^{i}_{t},{\bm{c}})}{p_{\theta_{\text{old}}}({\bm{x}}^{i}_{t-1}\mid{\bm{x}}^{i}_{t},{\bm{c}})}.$ (6)
-4.2 ODE-to-SDE
-In Eq.
+$f(r,\hat{A},\theta,\varepsilon,\beta)$ $=\frac{1}{G}\sum_{i=1}^{G}\frac{1}{T}\sum_{t=0}^{T-1}\Bigg{(}\min\Big{(}r^{i}_{t}(\theta)\hat{A}^{i}_{t},\ \text{clip}\Big{(}r^{i}_{t}(\theta),1-\varepsilon,1+\varepsilon\Big{)}\hat{A}^{i}_{t}\Big{)}-\beta D_{\text{KL}}(\pi_{\theta}||\pi_{\text{ref}})\Bigg{)},$
+$r^{i}_{t}(\theta)$ $=\frac{p_{\theta}({\bm{x}}^{i}_{t-1}\mid{\bm{x}}^{i}_{t},{\bm{c}})}{p_{\theta_{\text{old}}}({\bm{x}}^{i}_{t-1}\mid{\bm{x}}^{i}_{t},{\bm{c}})}.$
+From ODE to SDE.
+GRPO relies on stochastic sampling in Eq.
 4 and Eq.
-6, GRPO relies on a stochastic sampling procedure in order to generate a diverse batch of trajectories for advantage estimation and policy exploration.
-In the case of diffusion models, the forward process consists of a sequence of random Gaussian noising steps, and its reverse-time sampler is implemented via a Markov chain of Gaussians with progressively decreasing variance—equivalently, an approximate solver for the score-based SDE.
-Thus, diffusion sampling is naturally stochastic.
-While in the case of probability flow model, as previously discussed in Eq. 1, its forward process can be formulated as a deterministic ODE:
+5 to generate diverse trajectories for advantage estimation and exploration.
+Diffusion models naturally support this: the forward process adds Gaussian noise step by step, and the reverse process approximates a score-based SDE solver via a Markov chain with decreasing variance.
+In contrast, flow matching models use a deterministic ODE for the forward process:
 $\mathrm{d}{\bm{x}}_{t}={\bm{v}}_{t}\mathrm{d}t,$ (6)
-where ${\bm{v}}_{t}$ is modeled via the flow matching objective in Eq.
-2.
-A common sampling method is to discretize this ODE (e.g., via the explicit Euler method or a higher-order integrator) yielding a one-to-one mapping between successive time steps.
-For example, with a fixed step size $\Delta t$, the Euler scheme gives:
-${\bm{x}}_{t-1}={\bm{x}}_{t}+\Delta t{\bm{v}}_{\theta}({\bm{x}}_{t},t,{\bm{c}})$ (8)
-which produces a single, deterministic trajectory $\{{\bm{x}}_{T},{\bm{x}}_{T-1},\cdots,{\bm{x}}_{0}\}$ given any initial ${\bm{x}}_{T}$ and prompt ${\bm{c}}$.
-This deterministic approach fails to meet GRPO’s policy update requirements in two key ways: (1) Eq.
-6 requires computing $p({\bm{x}}_{t-1}\mid{\bm{x}}_{t},{\bm{c}})$, but deterministic sampling cannot provide probabilities.
-This limitation affects not only GRPO but all policy gradient methods, including the basic online REINFORCE [55] policy gradient, which also require probability computation.
-(2) Exploration is crucial for RL performance.
-As shown in Section 5.3, reduced randomness significantly lowers training efficiency.
-Deterministic sampling, which lacks randomness beyond the initial seed, is especially problematic.
+where ${\bm{v}}_{t}$ is learned via the flow matching objective in Eq. 2. A common sampling method is to discretize this ODE, yielding a one-to-one mapping between successive time steps.
+This deterministic approach fails to meet the GRPO policy update requirements in two key ways: (1) $r^{i}_{t}(\theta)$ in Eq.
+5 requires computing $p({\bm{x}}_{t-1}\mid{\bm{x}}_{t},{\bm{c}})$, which becomes computationally expensive under deterministic dynamics due to divergence estimation.
+(2) More importantly, RL depends on exploration.
+As shown in Section 5.3, reduced randomness greatly lowers training efficiency.
+Deterministic sampling, with no randomness beyond the initial seed, is especially problematic.
 To address this limitation, we convert the deterministic Flow-ODE from Eq.
@@
 A detailed proof is provided in Appendix A.
-Following [23, 28, 29], we construct a reverse-SDE formulation that preserves the marginal distribution:
+Following [23, 28, 29], we construct a reverse-time SDE formulation that preserves the marginal distribution:
 $\mathrm{d}{\bm{x}}_{t}=\left({\bm{v}}_{t}({\bm{x}}_{t})-\frac{\sigma_{t}^{2}}{2}\nabla\log p_{t}({\bm{x}}_{t})\right)\mathrm{d}t+\sigma_{t}\mathrm{d}{\bm{w}},$ (7)
-where $\mathrm{d}{\bm{w}}$ denotes Wiener process increments and $\sigma_{t}$ control the level of stachasticity during generation. For rectified flow, the marginal score has a relation with velocity:
-$\nabla\log p_{t}({\bm{x}})=-\frac{{\bm{x}}}{t}-\frac{1-t}{t}{\bm{v}}_{t}({\bm{x}}).$ (10)
-Substituting Eq. 10 into Eq. 9 gives the final SDE:
+where $\mathrm{d}{\bm{w}}$ denotes Wiener process increments and $\sigma_{t}$ control the level of stachasticity during generation. For rectified flow, Eq. 7 is specified as:
 $\mathrm{d}{\bm{x}}_{t}=\left[{\bm{v}}_{t}({\bm{x}}_{t})+\frac{\sigma_{t}^{2}}{2t}\left({\bm{x}}_{t}+(1-t){\bm{v}}_{t}({\bm{x}}_{t})\right)\right]\mathrm{d}t+\sigma_{t}\mathrm{d}{\bm{w}}.$ (8)
@@
 9 reveals that the policy $\pi_{\theta}({\bm{x}}_{t-1}\mid{\bm{x}}_{t},{\bm{c}})$ is an isotropic Gaussian distribution.
-We can easily compute the KL divergence between $\pi_{\theta}$ and the reference policy $\pi_{\text{ref}}$ as a closed form:
-$D_{\text{KL}}(\pi_{\theta}||\pi_{\text{ref}})=\frac{\|\overline{{\bm{x}}}_{t+\Delta t,\theta}-\overline{{\bm{x}}}_{t+\Delta t,\text{ref}}\|^{2}}{2\sigma_{t}^{2}\Delta t}=\frac{\Delta t}{2}\left(\frac{\sigma_{t}(1-t)}{2t}+\frac{1}{\sigma_{t}}\right)^{2}\|{\bm{v}}_{\theta}({\bm{x}}_{t},t)-{\bm{v}}_{\text{ref}}({\bm{x}}_{t},t)\|^{2}$ (13)
-4.3 Denoising Reduction
+We can easily compute the KL divergence between $\pi_{\theta}$ and the reference policy $\pi_{\text{ref}}$ in Eq.
+5 as a closed form:
+$D_{\text{KL}}(\pi_{\theta}||\pi_{\text{ref}})=\frac{\|\overline{{\bm{x}}}_{t+\Delta t,\theta}-\overline{{\bm{x}}}_{t+\Delta t,\text{ref}}\|^{2}}{2\sigma_{t}^{2}\Delta t}=\frac{\Delta t}{2}\left(\frac{\sigma_{t}(1-t)}{2t}+\frac{1}{\sigma_{t}}\right)^{2}\|{\bm{v}}_{\theta}({\bm{x}}_{t},t)-{\bm{v}}_{\text{ref}}({\bm{x}}_{t},t)\|^{2}$
+Denoising Reduction.
 To produce high-quality images, flow models typically require many denoising steps, making data collection costly for online RL.
@@
 5 Experiments
-In this section, we empirically evaluate Flow-GRPO’s ability to improving flow matching models on three representative tasks.
+This section empirically evaluates Flow-GRPO’s ability to improve flow matching models on three tasks.
 (1) Composition Image Generation: This task requires precise object arrangement and attribute control.
 We report the results on GenEval.
-(2) Visual Text Rendering: a rule-based task that stresses the accurate rendering of the text specified in the prompt.
+(2) Visual Text Rendering: a rule-based task that evaluates the accurate rendering of the text specified in the prompt.
 (3) Human Preference Alignment: This task aims to align T2I models with human preferences.
-We use PickScore [19] as our reward model.
 5.1 Experimental Setup
-We introduce three tasks, detailing their respective prompts and reward definitions. For hyperparameter details and compute resource specifications, please refer to Appendix B.2 and Appendix B.2.1.
-Compositional Image Generation
+We introduce three tasks, detailing their respective prompts and reward definitions. For hyperparameter details and compute resource specifications, please refer to Appendix B.3 and Appendix B.4.
+Compositional Image Generation.
 […]
-  • Counting: $r=|N_{\text{gen}}-N_{\text{ref}}|/N_{\text{ref}}$.
-  • Position / Color: If the object count is correct, a partial reward is assigned; the remainder is granted when the predicted position or color is also correct.
+Visual Text Rendering [8].
+Text is common in images such as posters, book covers, and memes, so the ability to place accurate and coherent text inside the generated images is crucial for T2I models.
+In our settings, we define an text rendering task, where each prompt follows the template ‘‘A sign that says ‘‘text’’.
+Specifically, the placeholder ‘‘text’’ is the exact string that should appear in the image.
+We use GPT4o to produce 20K training prompts and 1K test prompts.
+Following [56], we measure text fidelity with the reward $r=\max(1-N_{\text{e}}/N_{\text{ref}},0),$ where $N_{\text{e}}$ is the minimum edit distance between the rendered text and the target text and $N_{\text{ref}}$ is the number of characters inside the quotation marks in the prompt.
+This reward also serves as our metric of text accuracy.
+Human Preference Alignment [19].
+This task aims to align T2I models with human preferences.
+We use PickScore [19] as our reward model, which is based on large-scale human annotated pairwise comparisons of images generated from the same prompt.
+For each image and prompt pair, PickScore provides an overall score that evaluates multiple criteria, such as the alignment of the image with the prompt and its visual quality.
+Image Quality Evaluation Metric.
+Since the T2I model is trained to maximize a predefined reward, it is vulnerable to reward hacking, where the reward increases but image quality or diversity declines.
+This study aims to make online RL effective for T2I generation without noticeably compromising quality or diversity.
+To detect reward hacking beyond task-specific accuracy, we evaluate four automatic image quality metrics: Aesthetic Score [57], DeQA [58], ImageReward [32], and UnifiedReward [59] (see Appendix B.1 for details).
+All metrics are computed on DrawBench [1], a comprehensive benchmark with diverse prompts for T2I models.
 Table 1: GenEval Result. Best scores are in blue, second-best in green. Results for models other than SD3.5-M are from [7] or their original papers. Obj.: Object; Attr.: Attribution.
@@
 Flow-GRPO (w/ KL) 23.31 5.92 4.22 1.28 3.66
-Visual Text Rendering [8]
-Text is common in images such as posters, book covers, and memes, so the ability to place accurate and coherent text inside the generated images is crucial for T2V models.
-In our settings, we define an OCR task, where each prompt follows the template ‘‘A sign that says ‘‘text’’.
-Specifically, the placeholder ‘‘text’’ is the exact string that should appear in the image.
-We use GPT to produce 20K training prompts and 1K test prompts.
-Following [65], we measure text fidelity with the reward
-$r=\frac{|N_{\text{gen}}-N_{\text{ref}}|}{N_{\text{ref}}},$
-where $N_{\text{gen}}$ is the number of characters detected in the image and $N_{\text{ref}}$ is the number of characters inside the quotation marks in the prompt.
-This reward also serves as our metric of text accuracy.
-Human Preference Alignment [19]
-This task aims to align T2I models with human preferences.
-We use PickScore  [^2] https://huggingface.co/yuvalkirstain/PickScore_v1 as our reward model, which is based on large-scale human annotated pairwise comparisons of images generated from the same prompt.
-For each image and prompt pair, PickScore provides an overall score that evaluates multiple criteria, such as the alignment of the image with the prompt and its visual quality.
-Image Quality Evaluation Metric
-Because the T2I model is trained to maximise a predefined reward signal in RL, it is susceptible to reward hacking—situations in which the numerical reward rises but the quality or diversity of the generated images declines.
-This study aims to make online RL truly effective for T2I generation without compromising quality or diversity.
-To detect potential reward hacking beyond task-specific accuracy, we evaluate four automatic image quality metrics:
-  • Aesthetic score [66]: a CLIP-based linear regressor that predicts an image’s aesthetic score.
-  • DeQA score [67]: a multimodal large language model based image-quality assessment (IQA) model that quantifies how distortions, texture damage, and other low-level artefacts affect perceived quality.
-  • ImageReward [32]: a general purpose T2I human preference reward model that captures text–image alignment, visual fidelity, and harmlessness.
-  • UnifiedReward [68]: a recently proposed unified reward model for multimodal understanding and generation that currently achieves state-of-the-art performance on the human preference assessment leaderboard.
-All metrics are evaluated on DrawBench [1], a comprehensive and challenging benchmark with diverse prompts for T2I models.
 5.2 Main Results
@@
 Flow-GRPO improved text rendering ability, again without decreasing image quality metrics and preference scores on DrawBench.
-See Figures 8 and 9 in Appendix C.3 for related qualitative examples.
+See Figures 9, 10 & 11 in Appendix C.4 for related qualitative examples.
 For the Human Preference task, image quality did not decrease without KL regularization.
-However, we found that omitting KL caused a collapse in visual diversity—a form of reward hacking discussed further in Section 5.3.
+However, we found that omitting KL caused a collapse in visual diversity, a form of reward hacking discussed further in Section 5.3.
 These results demonstrate that Flow-GRPO boosts desired capabilities while causing very little degradation to image quality or visual diversity.
+We also compare Flow-GRPO with other alignment methods, including supervised fine-tuning, reward weighted regression [14, 69], Flow-DPO [14], and their online variants.
+Flow-GRPO outperforms all of them by a large margin.
+See Appendix C.1 for details.
 5.3 Analysis
-Reward Hacking
-We explored two approaches to mitigate reward hacking: (1) Multi-reward ensemble: We combined preference-based rewards (e.g., UnifiedReward, PickScore, ImageReward), image quality rewards (e.g., Aesthetic Score, DeQA) with task reward.
-However, this also led to local blurriness and reduced diversity.
-Due to the challenge of covering all aspects of generation with diverse reward models, we abandoned this approach.
-(2) KL constraint: We tuned the KL coefficient to maintain a small, constant KL divergence during training, keeping the model close to its pretrained weight.
-This allows for task-specific reward optimization without harming general performance.
-As shown in Table 2, removing the KL constraint boosts performance on Compositional Image Generation and Visual Text Rendering but causes a sharp decline in image quality and preference scores on DrawBench.
-In contrast, a properly tuned KL constraint preserves quality while maintaining similar gains on task-specific metrics.
-In the Human Preference Alignment task, dropping the KL constraint doesn’t hurt image quality—likely due to PickScore overlap with evaluation metrics—but causes a collapse in visual diversity.
+Reward Hacking.
+We use KL regularization to mitigate reward hacking by tuning the KL coefficient to keep the divergence small and nearly constant during training, keeping the model close to its pretrained weights.
+This allows task-specific reward optimization without harming overall performance.
+As shown in Table 2, removing the KL constraint for Compositional Image Generation and Visual Text Rendering significantly reduces image quality and preference scores on DrawBench.
+In contrast, a properly tuned KL preserves quality while achieving similar gains on task-specific metrics.
+In the Human Preference Alignment task, removing KL does not affect image quality, likely due to overlap between PickScore and evaluation metrics, but causes a collapse in visual diversity.
 Outputs converge to a single style, with different seeds producing nearly identical results.
-KL regularization prevents this collapse and preserves diversity.
-See Figure 7 in the Appendix C.2 for learning curves with and without KL, and Figure 4 for examples.
-Figure 4: Effect of KL Regularization. The KL penalty effectively suppresses reward hacking, preventing potential Quality Degradation (for GenEval and OCR) and Diversity Decline (for PickScore).
-Effect of Denoising Reduction
+KL regularization prevents this collapse and maintains diversity.
+See Figure 8 in Appendix C.3 for training curves and Figure 4 for more examples.
+Figure 4: Effect of KL Regularization. The KL penalty effectively suppresses reward hacking, preventing Quality Degradation (for GenEval and OCR) and Diversity Decline (for PickScore).
+Effect of Denoising Reduction.
 Figure 5 (a) highlights Denoising Reduction’s significant impact on accelerating training.
-To explore how different timesteps affect optimization, these experiments were conducted without the KL constraint.
-Reducing data collection timesteps from $40$ to $10$ achieved over a 4×4$\times$4 × speedup across all three tasks, without impacting final reward.
-While further tests showed this reduction didn’t consistently improve speed (and occasionally slowed some tasks), we standardized on $10$ timesteps for subsequent experiments, prioritizing the maintained reward.
+To explore how different timesteps affect optimization, these experiments are conducted without the KL constraint.
+Reducing data collection timesteps from $40$ to $10$ achieves over a 4×4$\times$4 × speedup across all three tasks, without impacting final reward.
+Further reducing to $5$ does not consistently improve speed and sometimes slows training, so we choose $10$ timesteps for later experiments.
 For the other two tasks, learning curves of reward versus training time are presented in Figure 7 in the Appendix C.2.
-Effect of Noise Level
+Effect of Noise Level.
 Higher $\sigma_{t}$ in the SDE boosts image diversity and exploration, vital for RL training.
 We control this exploration with a noise level $a$ (Eq.
-12).
-Ablation studies (Figure 5 (b)) show $a$’s impact on performance.
-Small $a$ (e.g., 0.1) limits exploration and slows reward improvement.
-Increasing $a$ (up to 0.7) enhances exploration and accelerates reward gains.
-Beyond this, further increases (e.g., $a=0.7\rightarrow 1.0$) yield no extra benefit, as exploration is sufficient.
-However, too much noise degrades image quality, leading to zero reward and failed training.
-We advise using the highest noise level that maintains good image quality.
+9).
+Figure 5 (b) shows the impact of $a$ on performance.
+A small $a$ (e.g., 0.1) limits exploration and slows reward improvement.
+Increasing $a$ (up to 0.7) boosts exploration and speeds up reward gains.
+Beyond this point (e.g., from 0.7 to 1.0), further increases provide no additional benefit, as exploration is already sufficient.
+We also observe that injecting too much noise by further increasing $a$ degrades image quality, resulting in zero reward and failed training.
 (a). Effect of Denoising Reduction on GenEval
@@
 (b) Noise Level: Moderate noise level ($a=0.7$) maximises OCR accuracy, while too little noise hampers exploration.
-Generalization Analysis
+Generalization Analysis.
 Flow-GRPO demonstrates strong generalization on unseen scenarios from GenEval (Table 4).
@@
 It also effectively controls object count, generalizing from training on $2-4$ objects to generate $5-6$ objects.
-Furthermore, Table 3 shows Flow-GRPO achieves significant gains on T2I-CompBench++.
+Furthermore, Table 3 shows Flow-GRPO achieves significant gains on T2I-CompBench++ [6, 70].
 This comprehensive benchmark for open-world compositional T2I generation features object classes and relationships substantially different from our model’s GenEval-style training data.
@@
 FLUX.1 Dev [5] 0.7407 0.5718 0.6922 0.2863 0.3866 0.6185 0.3127
-SD3.5-M 0.7994 0.5669 0.7338 0.2850 0.3739 0.5927 0.3146
+SD3.5-M [4] 0.7994 0.5669 0.7338 0.2850 0.3739 0.5927 0.3146
 SD3.5-M+Flow-GRPO 0.8379 0.6130 0.7236 0.5447 0.4471 0.6752 0.3195
@@
 SD3.5-M+Flow-GRPO 0.90 1.00 0.94 0.86 0.97 0.84 0.77 0.48
+6 Conclusion
+We have presented Flow-GRPO, the first method to integrate online RL into flow matching models.
+By converting deterministic ODEs to SDEs and reducing denoising steps during training, Flow-GRPO enables efficient RL-based optimization without noticeably compromising image quality or diversity.
+Our method significantly improves performance on compositional generation, text rendering, and human preference alignment, with minimal reward hacking.
+Flow-GRPO offers a simple and general framework for applying online RL to flow-based generative models.
 Limitations & Future Work.
-While this work focuses on T2I tasks, Flow-GRPO’s potential applicability to video generation [25, 27] raises several questions for future research: (1) Reward Design for Video: Defining effective reward models for video generation is crucial.
-Simple heuristics like object detectors or trackers can promote physical realism and temporal consistency, but more sophisticated models are needed.
-(2) Balancing Multiple Rewards: Effective video generation requires optimizing multiple objectives (e.g., realism, smoothness, coherence).
-Balancing these often conflicting signals is challenging and demands careful tuning.
-(3) Scalability: Video generation is significantly more resource-intensive than T2I.
-Scaling Flow-GRPO for video thus demands more efficient data collection and training pipelines.
+Although this work focuses on T2I tasks, Flow-GRPO has potential for video generation [25, 27], raising several future directions: (1) Reward Design: Simple heuristics, such as using object detectors or trackers as rule-based rewards, can encourage physical realism and temporal consistency, but more advanced reward models are needed.
+(2) Balancing Multiple Rewards: Video generation requires optimizing multiple objectives, including realism, smoothness, and coherence.
+Balancing these competing goals remains challenging and demands careful tuning.
+(3) Scalability: Video generation is far more resource-intensive than T2I, so applying Flow-GRPO at scale requires more efficient data collection and training pipelines.
+Additionally, better methods for preventing reward hacking are worth exploring.
+While KL regularization helps significantly, it requires longer training and occasional reward hacking occurs for certain prompts.
 Appendix A Mathematical Derivations for Stochastic Sampling using Flow Models
@@
 We consider its stochastic counterpart.
-Inspired by the derivation from SDE to its probability flow ODE in SGMs [23], we aim to construct a SDE with specific drift and diffusion coefficients so that its marginal distribution matches that of Eq.
-14.
+Inspired by the derivation from SDE to its probability flow ODE in SGMs [23], we aim to construct a forward SDE with specific drift and diffusion coefficients so that its marginal distribution matches that of Eq.
+10.
 We begin with the generic form of SDE:
@@
 To ensure that the stochastic process shares the same marginal distribution as the ODE, we impose:
-$-\nabla\cdot[f_{\text{SDE}}p_{t}({\bm{x}})]+\frac{1}{2}\nabla^{2}[\sigma_{t}^{2}p_{t}({\bm{x}})]=-\nabla\cdot[{\bm{v}}_{t}p_{t}({\bm{x}})]$ (18)
+$-\nabla\cdot[f_{\text{SDE}}\,p_{t}({\bm{x}})]+\frac{1}{2}\nabla^{2}[\sigma_{t}^{2}p_{t}({\bm{x}})]=-\nabla\cdot[{\bm{v}}_{t}({\bm{x}}_{t},t)p_{t}({\bm{x}})]$ (14)
 Observing that
@@
 $=\sigma_{t}^{2}\nabla\cdot(p_{t}({\bm{x}})\nabla\log p_{t}({\bm{x}}))$
-Substituting Eq. 19 to Eq. 18, we arrive at the drift coefficients of the target SDE:
-$f_{\text{SDE}}={\bm{v}}_{t}-\frac{1}{2}\sigma_{t}^{2}\nabla\log p_{t}({\bm{x}})$ (20)
-Hence we can rewrite the reverse-time SDE in Eq. 15 as:
+Substituting Eq. 15 to Eq. 14, we arrive at the drift coefficients of the target forward SDE:
+$f_{\text{SDE}}={\bm{v}}_{t}({\bm{x}}_{t},t)+\frac{\sigma_{t}^{2}}{2}\nabla\log p_{t}({\bm{x}})$ (16)
+Hence, we can rewrite the forward SDE in Eq. 11 as:
+$\mathrm{d}{\bm{x}}_{t}=\left({\bm{v}}_{t}({\bm{x}}_{t})+\frac{\sigma_{t}^{2}}{2}\nabla\log p_{t}({\bm{x}}_{t})\right)\mathrm{d}t+\sigma_{t}\mathrm{d}{\bm{w}},$ (17)
+where $\mathrm{d}{\bm{w}}$ denotes Wiener process increments, and $\sigma_{t}$ is the diffusion coefficient controlling the level of stochasticity during sampling.
+The relationship between forward and reverse-time SDEs has been established in [72, 23]. Specifically, if the forward SDE takes the form
+$\mathrm{d}{\bm{x}}_{t}=f({\bm{x}}_{t},t)\,\mathrm{d}t+g(t)\,\mathrm{d}{\bm{w}},$ (18)
+then the corresponding reverse-time SDE is
+$\mathrm{d}{\bm{x}}_{t}=\left[f({\bm{x}}_{t},t)-g^{2}(t)\nabla\log p_{t}({\bm{x}}_{t})\right]\,\mathrm{d}t+g(t)\,\mathrm{d}\overline{{\bm{w}}}.$ (19)
+Setting $g(t)=\sigma_{t}$, we obtain the reverse-time SDE corresponding to Eq. 17 as
+$\mathrm{d}{\bm{x}}_{t}$ $=\left[{\bm{v}}_{t}({\bm{x}}_{t})+\frac{\sigma_{t}^{2}}{2}\nabla\log p_{t}({\bm{x}}_{t})-\sigma_{t}^{2}\nabla\log p_{t}({\bm{x}}_{t})\right]\mathrm{d}t+\sigma_{t}\mathrm{d}\overline{{\bm{w}}}.$ (20)
+We thus arrive at the final form of the reverse-time SDE:
 $\mathrm{d}{\bm{x}}_{t}=\left({\bm{v}}_{t}({\bm{x}}_{t})-\frac{\sigma_{t}^{2}}{2}\nabla\log p_{t}({\bm{x}}_{t})\right)\mathrm{d}t+\sigma_{t}\mathrm{d}{\bm{w}},$ (21)
-where $d{\bm{w}}$ denotes Wiener process increments and $\sigma_{t}$ is the diffusion coefficient that controls the level of stachasticity during sampling.
-Once the score $\nabla\log p_{t}({\bm{x}}_{t})$ is available, the stochastic process can be sampled directly.
-While the Flow-Matching framework, this score is tied to the velocity field ${\bm{v}}_{t}$.
+Once the score function $\nabla\log p_{t}({\bm{x}}_{t})$ is available, the process can be simulated directly. For flow matching, this score is implicitly linked to the velocity field ${\bm{v}}_{t}$.
 Specifically, let $\dot{\alpha_{t}}\equiv\partial\alpha_{t}/\partial t$.
@@
 Appendix B Further Details on the Experimental Setup
-B.1 Model Specification
+B.1 Quality Metrics
+The details of quality metrics are as follows:
+  • Aesthetic score [57]: a CLIP-based linear regressor that predicts an image’s aesthetic score.
+  • DeQA score [58]: a multimodal large language model based image-quality assessment (IQA) model that quantifies how distortions, texture damage, and other low-level artefacts affect perceived quality.
+  • ImageReward [32]: a general purpose T2I human preference reward model that captures text–image alignment, visual fidelity, and harmlessness.
+  • UnifiedReward [59]: a recently proposed unified reward model for multimodal understanding and generation that currently achieves state-of-the-art performance on the human preference assessment leaderboard.
+B.2 Model Specification
 The following table lists the base model and the reward models and their corresponding links.
-B.2 Hyperparameters Specification
+B.3 Hyperparameters Specification
 Except for $\beta$, GRPO hyperparameters are fixed across tasks.
@@
 We use Lora with $\alpha=64$ and $r=32$.
-B.2.1 Compute Resources Specification
+B.4 Compute Resources Specification
 We train our model using 24 NVIDIA A800 GPUs. The learning curves in Appendix C.3 provide details on the specific GPU hours.
 Appendix C Extended Experimental Results
-C.1 Effect of Denoising Reduction
+C.1 Flow-GRPO vs. Other Alignment Methods
+We compare Flow-GRPO with several alignment methods: supervised fine-tuning (SFT), reward-weighted regression (Flow-RWR [14, 69]), Flow-DPO [14], and their online variants.
+Flow-GRPO consistently outperforms all baselines by a significant margin.
+At each step, we generate a group of images using the same group size as in Flow-GRPO.
+The only difference lies in the update rule:
+  • SFT: Select the highest-reward image in each group and fine-tune on it.
+  • Flow-RWR[14, 69]: Apply a softmax over rewards in each group and perform reward-weighted likelihood maximization.
+  • Flow-DPO[14]: Use the highest-reward image in each group as the chosen sample and the lowest as the rejected, then apply the DPO loss.
+Offline variants use a fixed pretrained model for data collection, while online variants update their data collection model every 40 steps.
+As shown in Figure 6, Flow-GRPO outperforms all other methods.
+The figure also indicates that DPO and SFT improve over time.
+In contrast, RWR does not, which aligns with experimental findings on RWR in [12].
+Additionally, Online DPO surpasses offline DPO, aligning with [15]’s finding that online DPO performs better.
+For the second-best online DPO, a hyperparameter search on its key parameter $\beta$ revealed that smaller values are not always optimal; excessively small $\beta$ values can cause training collapse.
+Figure 6: Comparison of Flow-GRPO and Other Alignment Methods.
+C.2 Effect of Denoising Reduction
 We show the extended Denoising Reduction ablations of Visual Text Rendering and Human Preference Alignment tasks in Figure 7.
@@
 Figure 7: Effect of Denoising Reduction
-C.2 Learning Curves with or without KL
+C.3 Learning Curves with or without KL
 Figure 8 shows learning curves for three tasks, with and without KL.
@@
 Figure 8: Learning Curves with and without KL. KL penalty slows early training yet effectively suppresses reward hacking.
-C.3 Qualitative Results
-In addition to the GenEval reward comparisons in Figure 3, Figures 8 and 9 show qualitative comparisons between SD-3.5-M and SD-3.5-M + Flow-GRPO trained with PickScore and OCR rewards, respectively.
-Figure 8: Qualitative comparison between the SD-3.5-M and SD-3.5-M + Flow-GRPO trained with PickScore reward.
-Figure 9: Qualitative comparison between the SD-3.5-M and SD-3.5-M + Flow-GRPO trained with OCR reward.
+C.4 Additional Qualitative Results
+Figures 9, 10 & 11 qualitatively compare SD3.5-M with its Flow-GRPO enhanced versions (with and without KL regularization) using GenEval, OCR and PickScore rewards, respectively.
+Flow-GRPO with KL regularization improves the target capability while maintaining image quality and minimizing reward-hacking.
+Conversely, removing the KL constraint significantly degrades image quality and diversity.
+Figure 9: Additional Qualitative comparison between the SD3.5-M and SD3.5-M + Flow-GRPO trained with GenEval reward.
+Figure 10: Additional Qualitative comparison between the SD3.5-M and SD3.5-M + Flow-GRPO trained with OCR reward.
+Figure 11: Additional Qualitative comparison between the SD3.5-M and SD3.5-M + Flow-GRPO trained with PickScore reward.
+C.5 Evolution of Evaluation Images During Flow-GRPO Training
+To better understand the training dynamics of our proposed Flow-GRPO framework, we visualize the evolution of generated samples corresponding to fixed evaluation prompts at regular intervals during training in Figure 121314.
+For consistency, all visualizations are produced using a 40-step ODE-based sampling schedule.
+These qualitative results provide a visual representation of how the model progressively improves its generation quality and alignment with task objectives over time.
+Figure 12: We visualize the generated samples across successive training iterations during the optimization of SD3.5-Medium on the GenEval task.
+Figure 13: We visualize the generated samples across successive training iterations during the optimization of SD3.5-Medium on the OCR task.
+Figure 14: We visualize the generated samples across successive training iterations during the optimization of SD3.5-Medium on the PickScore task.
+Appendix D Training Sample Visualization with Denoising Reduction
+In this section, we compare images obtained with SDE sampling at various steps against those produced by ODE sampling, and offer an intuitive view of the denoising reduction strategy.
+Figure 15 presents SD3.5-Medium samples under four inference settings: (a) ODE sampling with 40 steps; (b) SDE sampling with 40 steps; (c) SDE sampling with 10 steps; (d) SDE sampling with 5 steps.
+The 40-step ODE and SDE runs yield visually indistinguishable images, confirming that our SDE sampler preserves quality.
+Shortening the SDE schedule to 10 and 5 steps introduces conspicuous artifacts, like color drift and fine details blur.
+Contrary to expectation that such low-quality samples might hinder optimization.
+it actually do just the opposite and accelerate optimization.
+Because Flow-GRPO relies on relative preferences, it still extracts a useful reward signal, while the shorter trajectories signifactly cut wall-clock time.
+Consequently, Flow-GRPO with denoising reduction strategy converges more quickly on both layout-oriented benchmarks such as GenEval and quality-focused metrics such as PickScore, without sacrificing final performance.
+Figure 15: Visualization of training samples under difference inference settings.

### v4

@@
 Figure 3: Qualitative Comparison on the GenEval Benchmark. Our approach demonstrates superior performance in Counting, Colors, Attribute Binding, and Position.
-Table 2: Performance on Compositional Image Generation, Visual Text Rendering, and Human Preference benchmarks, evaluated by task performance, image quality, and preference scores.
-Acc: Accuracy; ImgRwd: ImageReward; UniRwd: UnifiedReward.
+Table 2: Performance on Compositional Image Generation, Visual Text Rendering, and Human Preference benchmarks, evaluated by task performance on test prompts, and by image quality and preference scores on DrawBench prompts.
+ImgRwd: ImageReward; UniRwd: UnifiedReward.
 Model Task Metric Image Quality Preference Score
@@
 Human Preference Alignment
-Flow-GRPO (w/o KL) 23.41 6.15 4.16 1.24 3.57
-Flow-GRPO (w/ KL) 23.31 5.92 4.22 1.28 3.66
+Flow-GRPO (w/o KL) 23.41 6.15 4.16 1.24 23.56 3.57
+Flow-GRPO (w/ KL) 23.31 5.92 4.22 1.28 23.53 3.66
 5.2 Main Results
@@
 Flow-GRPO improved text rendering ability, again without decreasing image quality metrics and preference scores on DrawBench.
-See Figures 9, 10 & 11 in Appendix C.4 for related qualitative examples.
+See Figures 10, 11 & 12 in Appendix C.4 for related qualitative examples.
 For the Human Preference task, image quality did not decrease without KL regularization.
@@
 While KL regularization helps significantly, it requires longer training and occasional reward hacking occurs for certain prompts.
+Acknowledgements
+We would like to thank Mingwu Zheng for discussions on the proof and Zhanhui Zhou for suggestions on the writing.
 Appendix A Mathematical Derivations for Stochastic Sampling using Flow Models
@@
 Other settings include a group size $G=24$, an noise level $a=0.7$ and an image resolution of 512.
-The KL ratio $\beta$ is set to 0.004 for GenEval and Text Rendering, and 0.001 for Pickscore.
+The KL ratio $\beta$ is set to 0.04 for GenEval and Text Rendering, and 0.01 for Pickscore.
 We use Lora with $\alpha=64$ and $r=32$.
@@
 For the second-best online DPO, a hyperparameter search on its key parameter $\beta$ revealed that smaller values are not always optimal; excessively small $\beta$ values can cause training collapse.
-Figure 6: Comparison of Flow-GRPO and Other Alignment Methods.
+Figure 7 shows similar conclusions.
+Figure 6: Comparison of Flow-GRPO and Other Alignment Methods on the Human Preference Alignment task.
+Figure 7: Comparison of Flow-GRPO and Other Alignment Methods on the Compositional Image Generation task.
 C.2 Effect of Denoising Reduction
 We show the extended Denoising Reduction ablations of Visual Text Rendering and Human Preference Alignment tasks in Figure 8.
-(a). Visual Text Rendering
-(b). Human Preference Alignment
+(a) Visual Text Rendering
+(b) Human Preference Alignment
 Figure 8: Effect of Denoising Reduction
@@
 Adding appropriate KL can achieve the same high reward as the KL-free version and maintain image quality, though it requires longer training.
-(a). Compositional Image Generation
-(b). Visual Text Rendering
-(c). Human Preference Alignment
+(a) Compositional Image Generation
+(b) Visual Text Rendering
+(c) Human Preference Alignment
 Figure 9: Learning Curves with and without KL. KL penalty slows early training yet effectively suppresses reward hacking.
 C.4 Additional Qualitative Results
-Figures 9, 10 & 11 qualitatively compare SD3.5-M with its Flow-GRPO enhanced versions (with and without KL regularization) using GenEval, OCR and PickScore rewards, respectively.
+Figures 10, 11 & 12 qualitatively compare SD3.5-M with its Flow-GRPO enhanced versions (with and without KL regularization) using GenEval, OCR and PickScore rewards, respectively.
 Flow-GRPO with KL regularization improves the target capability while maintaining image quality and minimizing reward-hacking.

### v5

@@
 Abstract
-We propose Flow-GRPO, the first method to integrate online reinforcement learning (RL) into flow matching models.
+We propose Flow-GRPO, the first method to integrate online policy gradient reinforcement learning (RL) into flow matching models.
 Our approach uses two key strategies: (1) an ODE-to-SDE conversion that transforms a deterministic Ordinary Differential Equation (ODE) into an equivalent Stochastic Differential Equation (SDE) that matches the original model’s marginal distribution at all timesteps, enabling statistical sampling for RL exploration; and (2) a Denoising Reduction strategy that reduces training denoising steps while retaining the original number of inference steps, significantly improving sampling efficiency without sacrificing performance.
@@
 Notably, very little reward hacking occurred, meaning rewards did not increase at the cost of appreciable image quality or diversity degradation.
-Figure 1: (a) GenEval performance rises steadily throughout Flow-GRPO’s training and outperforms GPT-4o.
+Figure 1: (a) GenEval performance rises steadily throughout Flow-GRPO’s training and outperforms GPT‑4o.
 (b) Image quality metrics on DrawBench [1] remain essentially unchanged.
@@
 RL for LLM.
-Online RL has effectively improved the reasoning abilities of LLMs, such as DeepSeek-R1 [10] and OpenAI-o1 [11], using policy gradient methods like PPO [20] or value-free GRPO [16].
+Online RL has effectively improved the reasoning abilities of LLMs, such as DeepSeek‑R1 [10] and OpenAI‑o1 [11], using policy gradient methods like PPO [20] or value-free GRPO [16].
 GRPO is more memory efficient by removing the need for a value network, so we adopt it in this work.
@@
 Alignment for T2I.
-Recent efforts to align pretrained T2I models with human preferences follow five main directions: (1) direct fine-tuning with scalar rewards [30, 31, 32, 33]; (2) Reward Weighted Regression (RWR) [34, 35, 36]; (3) Direct Preference Optimization (DPO) and variants [37, 38, 39, 14, 40, 41, 42, 43, 44, 36]; (4) PPO-style policy gradients [45, 46, 47, 48, 49, 50]; (5) training-free alignment methods [51, 52, 53].
+Recent efforts to align pretrained T2I models with human preferences follow five main directions: (1) direct fine-tuning with differentiable rewards [30, 31, 32, 33]; (2) Reward Weighted Regression (RWR) [34, 35, 36, 37]; (3) Direct Preference Optimization (DPO) and variants [38, 39, 14, 40, 41, 42, 43, 44, 45, 46]; (4) PPO-style policy gradients [47, 48, 49, 50, 51, 52]; (5) training-free alignment methods [53, 54, 55].
 These methods have successfully aligned T2I models with human preferences, improving aesthetics and semantic consistency.
@@
 We report the results on GenEval.
-(2) Visual Text Rendering: a rule-based task that evaluates the accurate rendering of the text specified in the prompt.
+(2) Visual Text Rendering: a rule‑based task that evaluates the accurate rendering of the text specified in the prompt.
 (3) Human Preference Alignment: This task aims to align T2I models with human preferences.
@@
 Compositional Image Generation.
-[…]
+GenEval [17] assesses T2I models on complex compositional prompts—like object counting, spatial relations, and attribute binding—across six difficult compositional image generation tasks.
+We use its official evaluation pipeline, which detects object bounding boxes and colors, then infers their spatial relations.
+Training prompts are generated using official GenEval scripts, which apply templates and random combinations to construct the prompt dataset.
+The test set is strictly deduplicated: prompts differing only in object order (e.g., "a photo of A and B" vs.
+"a photo of B and A") are treated as identical, and these variants are removed from the training set.
+Based on the base model’s initial accuracy across the six tasks, we set the prompt ratio as $\texttt{Position}:\texttt{Counting}:\texttt{Attribute Binding}:\texttt{Colors}:\texttt{Two Objects}:\texttt{Single Object}=7:5:3:1:1:0$.
+Rewards are rule-based: (1) Counting: $r=1-|N_{\text{gen}}-N_{\text{ref}}|/N_{\text{ref}}$; (2) Position / Color: If the object count is correct, a partial reward is assigned; the remainder is granted when the predicted position or color is also correct.
 Visual Text Rendering [8].
@@
 Flow-GRPO (w/ KL) 23.31 5.92 4.22 1.28 23.53 3.66
+Figure 4: Comparison with Other Alignment Methods on the Compositional Generation Task.
+Figure 5: Ablation Studies on Different Group Size $G$. Higher group size performs better.
 5.2 Main Results
@@
 Flow-GRPO improved text rendering ability, again without decreasing image quality metrics and preference scores on DrawBench.
-See Figures 10, 11 & 12 in Appendix C.4 for related qualitative examples.
+See Figures 13, 14 & 15 in Appendix C.6 for related qualitative examples.
 For the Human Preference task, image quality did not decrease without KL regularization.
@@
 These results demonstrate that Flow-GRPO boosts desired capabilities while causing very little degradation to image quality or visual diversity.
-We also compare Flow-GRPO with other alignment methods, including supervised fine-tuning, reward weighted regression [14, 69], Flow-DPO [14], and their online variants.
-Flow-GRPO outperforms all of them by a large margin.
-See Appendix C.1 for details.
+Flow-GRPO vs. Other Alignment Methods.
+We compare Flow-GRPO with several alignment methods: supervised fine-tuning (SFT), Flow-DPO [14, 39], and their online variants.
+Flow-GRPO consistently outperforms all baselines by a significant margin.
+At each step, we generate a group of images using the same group size as in Flow-GRPO.
+The only difference lies in the update rule:
+  • SFT: Select the highest-reward image in each group and fine-tune on it.
+  • Flow-DPO: Use the highest-reward image in each group as the chosen sample and the lowest as the rejected, then apply the DPO loss.
+Offline variants use a fixed pretrained model for data collection, while online variants update their data collection models every 40 steps.
+As shown in Figure 5, Flow-GRPO outperforms all baselines.
+Online DPO also surpasses its offline counterpart, consistent with [15].
+For the second-best online DPO, a hyperparameter search on its key parameter $\beta$ revealed that smaller values are not always optimal; excessively small $\beta$ values can cause training collapse.
+Appendix C presents more comprehensive comparisons covering additional methods and tasks.
 5.3 Analysis
+This section presents several analyses to better understand the behavior and robustness of Flow-GRPO.
+We examine issues such as reward hacking, the impact of denoising reduction and noise levels, the effect of group size, and the model’s generalization ability.
+We provide additional analyses in the Appendix C.
 Reward Hacking.
@@
 To explore how different timesteps affect optimization, these experiments are conducted without the KL constraint.
-Reducing data collection timesteps from $40$ to $10$ achieves over a 4×4$\times$4 × speedup across all three tasks, without impacting final reward.
+Reducing data collection timesteps from $40$ to $10$ achieves over a $4\times$ speedup across all three tasks, without impacting final reward.
 Further reducing to $5$ does not consistently improve speed and sometimes slows training, so we choose $10$ timesteps for later experiments.
@@
 (b) Noise Level: Moderate noise level ($a=0.7$) maximises OCR accuracy, while too little noise hampers exploration.
+Effect of Group Size.
+Figure 5 shows the effect of group size $G$ using PickScore as the reward function.
+When the group size was reduced to $G=12$ and $G=6$, training became unstable and eventually collapsed, whereas $G=24$ remained stable throughout the process.
+We observe that smaller group sizes produce inaccurate advantage estimates, increasing variance and leading to training collapse, a phenomenon also reported in [71, 72].
 Generalization Analysis.
@@
 Specifically, it captures object number, color, and spatial relations, generalizing well to unseen object classes.
-It also effectively controls object count, generalizing from training on $2-4$ objects to generate $5-6$ objects.
+It also effectively controls object count, generalizing from training on 242-4 objects to generate 565-6 or $12$ objects.
 Furthermore, Table 3 shows Flow-GRPO achieves significant gains on T2I-CompBench++ [6, 73].
@@
 Unseen Objects: Trained on 60 object classes, evaluated on 20 unseen classes.
-Unseen Counting: Trained to render 2, 3, or 4 objects, was evaluated on rendering 5 or 6.
+Unseen Counting: Trained to render 2, 3, or 4 objects, and evaluated in two settings: rendering 5 or 6 objects, and rendering 12 objects.
 Method Unseen Objects Unseen Counting
-Overall Single Obj. Two Obj. Counting Colors Position Attr. Binding Counting
-SD3.5-M 0.64 0.96 0.73 0.53 0.87 0.26 0.47 0.13
-SD3.5-M+Flow-GRPO 0.90 1.00 0.94 0.86 0.97 0.84 0.77 0.48
+Overall Single Obj. Two Obj. Counting Colors Position Attr. Binding 5-6 Objects 12 Objects
+SD3.5-M 0.64 0.96 0.73 0.53 0.87 0.26 0.47 0.13 0.02
+SD3.5-M+Flow-GRPO 0.90 1.00 0.94 0.86 0.97 0.84 0.77 0.48 0.12
 6 Conclusion
-We have presented Flow-GRPO, the first method to integrate online RL into flow matching models.
+We have presented Flow-GRPO, the first method to integrate online policy gradient RL into flow matching models.
 By converting deterministic ODEs to SDEs and reducing denoising steps during training, Flow-GRPO enables efficient RL-based optimization without noticeably compromising image quality or diversity.
@@
 Acknowledgements
-We would like to thank Mingwu Zheng for discussions on the proof and Zhanhui Zhou for suggestions on the writing.
+This work was partially supported by the JC STEM Lab of AI for Science and Engineering, funded by The Hong Kong Jockey Club Charities Trust, the Research Grants Council of Hong Kong (Project No.
+CUHK14213224).
+We gratefully acknowledge Mingwu Zheng for his insightful discussions on the proof and Zhanhui Zhou for his valuable comments that improved the clarity of this paper.
 Appendix A Mathematical Derivations for Stochastic Sampling using Flow Models
@@
 To ensure that the stochastic process shares the same marginal distribution as the ODE, we impose:
-$-\nabla\cdot[f_{\text{SDE}}\,p_{t}({\bm{x}})]+\frac{1}{2}\nabla^{2}[\sigma_{t}^{2}p_{t}({\bm{x}})]=-\nabla\cdot[{\bm{v}}_{t}({\bm{x}}_{t},t)p_{t}({\bm{x}})]$ (14)
+$\nabla\cdot[f_{\text{SDE}}\,p_{t}({\bm{x}})]+\frac{1}{2}\nabla^{2}[\sigma_{t}^{2}p_{t}({\bm{x}})]=-\nabla\cdot[{\bm{v}}_{t}({\bm{x}}_{t},t)p_{t}({\bm{x}})]$ (14)
 Observing that
@@
 The details of quality metrics are as follows:
-  • Aesthetic score [57]: a CLIP-based linear regressor that predicts an image’s aesthetic score.
-  • DeQA score [58]: a multimodal large language model based image-quality assessment (IQA) model that quantifies how distortions, texture damage, and other low-level artefacts affect perceived quality.
+  • Aesthetic score [59]: a CLIP‑based linear regressor that predicts an image’s aesthetic score.
+  • DeQA score [60]: a multimodal large language model based image‑quality assessment (IQA) model that quantifies how distortions, texture damage, and other low‑level artefacts affect perceived quality.
   • ImageReward [32]: a general purpose T2I human preference reward model that captures text–image alignment, visual fidelity, and harmlessness.
-  • UnifiedReward [59]: a recently proposed unified reward model for multimodal understanding and generation that currently achieves state-of-the-art performance on the human preference assessment leaderboard.
+  • UnifiedReward [61]: a recently proposed unified reward model for multimodal understanding and generation that currently achieves state‑of‑the‑art performance on the human preference assessment leaderboard.
 B.2 Model Specification
@@
   • Flow-RWR [14, 76]: Apply a softmax over rewards in each group and perform reward-weighted likelihood maximization.
-  • Flow-DPO[14]: Use the highest-reward image in each group as the chosen sample and the lowest as the rejected, then apply the DPO loss.
+  • Flow-DPO [14, 39]: Use the highest-reward image in each group as the chosen sample and the lowest as the rejected, then apply the DPO loss.
 Offline variants use a fixed pretrained model for data collection, while online variants update their data collection model every 40 steps.
@@
 For the second-best online DPO, a hyperparameter search on its key parameter $\beta$ revealed that smaller values are not always optimal; excessively small $\beta$ values can cause training collapse.
-Figure 7 shows similar conclusions.
 Figure 8: Comparison of Flow-GRPO and Other Alignment Methods on the Human Preference Alignment task.
-Figure 7: Comparison of Flow-GRPO and Other Alignment Methods on the Compositional Image Generation task.
+Since methods like DPO use different tuned batch sizes from Flow-GRPO, we use the number of training prompts on the x-axis for a fair comparison across these methods.
+DDPO.
+DDPO [12] was originally developed for diffusion-based backbones, so we adapted it to flow-matching models via our ODE-to-SDE conversion.
+Using SD3.5-M as the base model and PickScore as the reward signal, we track the evaluation reward throughout the entire training process in Figure 8.
+We find that DDPO’s reward increases more slowly than Flow-GRPO’s and eventually collapses in the later stages, whereas Flow-GRPO trains stably and continues to improve consistently over time.
+ReFL.
+ReFL [32] directly fine-tunes diffusion models by viewing reward model scores as human preference losses and back-propagating gradients to a randomly-picked late timestep $t$.
+Following ImageReward [32], we back-propagate gradients to a randomly chosen late timestep $t\in[30,40]$ during denoising.
+Figure 8 shows that GRPO surpasses ReFL when the reward is differentiable, indicating that GRPO maintains strong performance in settings where ReFL applies.
+More importantly, GRPO does not require differentiable rewards, enabling direct use of state-of-the-art Vision-Language Models (VLMs) as reward providers.
+This offers two key advantages:
+  • Sophisticated, General-Purpose Rewards: VLMs can conduct human-like evaluations through a structured reasoning process.
+Given a prompt, a VLM can decompose it into key criteria, reason step by step to verify each aspect in the generated image, and then provide a comprehensive overall score.
+This enables a single, unified reward model to handle diverse tasks, from text-to-image generation to complex instruction-based image editing.
+  • Future-Proof and Cost-Free Upgrades: The field of VLMs is advancing at a breathtaking pace.
+By using a VLM as the reward source, our framework automatically benefits from these improvements.
+As VLMs become more capable, the reward model becomes stronger without any additional training data or computational cost.
+ORW.
+ORW [35] is an online reward-weighted regression method that guides the model to prioritize high-reward regions.
+Unlike KL regularization, it employs Wasserstein-2 regularization to prevent policy collapse and maintain diversity.
+To ensure a fair comparison, we adopt the same experimental setup as in our Human Preference Alignment task.
+For ORW, we set $\beta=0.5$ and $\alpha=1$ (lower values led to unstable training).
+The steps_per_epoch parameter, which controls how frequently the data-collecting policy is updated, was chosen from 20,40,100,400{20,40,100,400} based on best performance.
+Table 5 reports reward scores on the test set across training steps.
+Following ORW’s Table 1, we randomly sampled 50 DrawBench prompts and generated 64 images per prompt to compute CLIP and Diversity scores.
+As shown in Table 6, Flow-GRPO outperforms ORW on both metrics.
+Table 5: Reward scores on the test set over training steps.
+Method Step 0 Step 240 Step 480 Step 720 Step 960
+SD3.5-M + ORW 28.79 29.05 29.15 27.58 23.05
+SD3.5-M + Flow-GRPO 28.79 29.10 29.17 29.51 29.89
+Table 6: Comparison of CLIP and diversity scores across different fine-tuning methods.
+   Method    CLIP Score $\uparrow$    Diversity Score $\uparrow$
+   SD3.5-M    27.99    0.96
+   SD3.5-M + ORW    28.40    0.97
+   SD3.5-M + Flow-GRPO    30.18    1.02
 C.2 Effect of Denoising Reduction
@@
 Figure 9: Effect of Denoising Reduction
-C.3 Learning Curves with or without KL
+C.3 Effect of Initial Noise
+We initialize each rollout with difference random noise to increase exploratory diversity during RL training.
+We perform an additioanl ablation to confirm this claim.
+With SD3.5-M as the base model and PickScore as the reward, we compare Flow-GRPO with different initial noise against Flow-GRPO with the same initial noise.
+Figure 11 shows the variant with different noise consistently achieved high rewards during the training process.
+Figure 10: Effect of Initial Noise
+Figure 11: Additional Results on FLUX.1-Dev
+C.4 Additional Results on FLUX.1-Dev
+We run Flow-GRPO on FLUX.1-Dev [5] using PickScore as the reward signal.
+The reward curve rises steadily throughout training without noticeable reward hacking.
+Figure 11 shows the reward values over the training process, and Table 7 compares FLUX.1-Dev with FLUX.1-Dev + Flow-GRPO on DrawBench.
+Table 7: Comparison of FLUX.1-Dev and Flow-GRPO fine-tuned models.
+Model Aesthetic DeQA ImageReward PickScore UnifiedReward
+FLUX.1-Dev 5.71 4.31 0.85 22.62 3.65
+FLUX.1-Dev + Flow-GRPO 6.02 4.24 1.32 23.97 3.81
+C.5 Learning Curves with or without KL
 Figure 12 shows learning curves for three tasks, with and without KL.
@@
 Figure 12: Learning Curves with and without KL. KL penalty slows early training yet effectively suppresses reward hacking.
-C.4 Additional Qualitative Results
-Figures 10, 11 & 12 qualitatively compare SD3.5-M with its Flow-GRPO enhanced versions (with and without KL regularization) using GenEval, OCR and PickScore rewards, respectively.
+C.6 Additional Qualitative Results
+Figures 13, 14 & 15 qualitatively compare SD3.5-M with its Flow-GRPO enhanced versions (with and without KL regularization) using GenEval, OCR and PickScore rewards, respectively.
 Flow-GRPO with KL regularization improves the target capability while maintaining image quality and minimizing reward-hacking.
@@
 Figure 15: Additional Qualitative comparison between the SD3.5-M and SD3.5-M + Flow-GRPO trained with PickScore reward.
-C.5 Evolution of Evaluation Images During Flow-GRPO Training
+C.7 Evolution of Evaluation Images During Flow-GRPO Training
 To better understand the training dynamics of our proposed Flow-GRPO framework, we visualize the evolution of generated samples corresponding to fixed evaluation prompts at regular intervals during training in Figure 161718.